### v9

# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

Abstract
[…]
1 Introduction
The multinomial logistic (MNL) bandit framework [45, 46, 6, 7, 38, 39, 42, 4, 51] describes sequential assortment selection problems in which an agent offer a sequence of assortments of at most $K$ items from a set of $N$ possible items and receives feedback only for the chosen decisions. The choice probability of each outcome is characterized by an MNL model [35]. This framework allows modeling of various real-world situations such as recommender systems and online retails, where selections of assortments are evaluated based on the user-choice feedback among offered multiple options.
In this paper, we study the contextual MNL bandit problem [7, 6, 41, 14, 38, 39, 42, 4], where the features of items and possibly contextual information about a user at each round are available. Despite many recent advances, [14, 38, 39, 42, 4], however, no previous studies have proven the minimax optimality of contextual MNL bandits. Chen et al. [14] proposed a regret lower bound of $\Omega(d\sqrt{T}/K)$, where $d$ is the feature dimension, $T$ is the total number of rounds, and $K$ is the maximum size of assortments, assuming the uniform rewards, i.e., rewards are all same for each of the total $N$ items. Furthermore, Chen and Wang [13] established a regret lower bound of $\Omega(\sqrt{NT})$ in the non-contextual setting (hence, dependence on $N$ appears instead of $d$), which is tighter in terms of $K$. It is important to note the difference in the assumptions for the attraction parameter for the outside option $v_{0}$. Chen and Wang [13] assumed for the attraction parameter for the outside option to be $v_{0}=K$, whereas Chen et al. [14] assumed $v_{0}=1$. Therefore, it remains an open question whether and how the value of $v_{0}$ affects both lower and upper bounds of regret.
Regarding regret upper bounds, Chen et al. [14] proposed an exponential runtime algorithm that achieves a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$ in the setting with stochastic contexts and the non-uniform rewards. Under the same setting, Oh and Iyengar [39] and Oh and Iyengar [38] introduced polynomial-time algorithms that attain regrets of $\tilde{\mathcal{O}}(d\sqrt{T}/\kappa)$ and $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T}/\kappa)$ respectively, where $1/\kappa=\mathcal{O}(K^{2})$ is a problem-dependent constant. Recently, Perivier and Goyal [42] improved the dependency on $\kappa$ in the adversarial context setting, achieving a regret of $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T})$, where $\kappa^{\prime}=\mathcal{O}(1/K)$. However, their approach focuses solely on the setting with uniform rewards, which is a special case of non-uniform rewards, and currently, there is no tractable method to implement the algorithm.
As summarized in Table 1, there has been a gap between the upper and lower bounds in the existing works of contextual MNL bandits. No previous studies have confirmed whether lower or upper bounds are tight, obscuring what the optimal regret should be. This ambiguity is further exacerbated because many studies introduce their methods under varying conditions such as different reward structures and values of $v_{0}$, without explicitly explaining how these factors impact regret. Additionally, there is currently no computationally efficient algorithm whose regret does not scale with $1/\kappa=\mathcal{O}(K^{2})$ or directly with $K$. Intuitively, increasing $K$ provides more information at least in the uniform reward setting, potentially leading to a more statistically efficient learning process. However, no previous results have reflected such intuition. Hence, the following research questions arise:
  • What is the optimal regret lower bound in contextual MNL bandits?
  • Can we design a computationally efficient, nearly minimax optimal algorithm under the adversarial context setting?
Table 1: Comparisons of lower and upper regret bounds in related works on MNL bandits with $T$ rounds, NN items, the maximum size of assortments $K$, $d$-dimensional feature vectors, and problem-dependent constants $1/\kappa=\mathcal{O}(K^{2})$ and $\kappa^{\prime}=\mathcal{O}(1/K)$. $\tilde{\mathcal{O}}$ represents big-$\mathcal{O}$ notation up to logarithmic factors. For the computational cost (abbreviated as “Comput.”), we consider only the dependence on the number of rounds $t$. “Intractable” means a non-polynomial runtime. The notation “-” denotes not applicable. The starred () papers only consider the non-contextual setting.
Regret Contexts Rewards $v_{0}$ Comput. per Round
Lower
Bound
Chen et al. [14] $\Omega(d\sqrt{T}/K)$ - Uniform $\Theta(1)$ -
Agrawal et al. [7] $\Omega(\sqrt{NT/K})$ - Uniform $\Theta(K)$ -
Chen and Wang [13] $\Omega(\sqrt{NT})$ - Uniform $\Theta(K)$ -
This work (Theorem 1) $\Omega(\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T})$ - Uniform Any value -
This work (Theorem 3) $\Omega(d\sqrt{T})$ - Non-uniform $\Theta(1)$ -
Upper
Bound
Chen et al. [14] $\tilde{\mathcal{O}}(d\sqrt{T})$ Stochastic Non-uniform $\Theta(1)$ Intractable
Oh and Iyengar [39] $\tilde{\mathcal{O}}(d\sqrt{T}/\kappa)$ Stochastic Non-uniform $\Theta(1)$ $\mathcal{O}(t)$
Oh and Iyengar [38] $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T}/\kappa)$ Adversarial Non-uniform $\Theta(1)$ $\mathcal{O}(t)$
Perivier and Goyal [42] $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T})$ Adversarial Uniform $\Theta(1)$ Intractable
This work (Theorem 2) $\tilde{\mathcal{O}}\left(\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T}\right)$ Adversarial Uniform Any value $\mathcal{O}(1)$
This work (Theorem 4) $\tilde{\mathcal{O}}(d\sqrt{T})$ Adversarial Non-uniform $\Theta(1)$ $\mathcal{O}(1)$
[…]
  • Under uniform rewards, we establish a regret lower bound of $\Omega(\sqrt{v_{0}K}/(v_{0}+K)d\sqrt{T})$ (Theorem 1), which is the tightest known lower bound in contextual MNL bandits. We propose, for the first time, a computationally efficient and provably optimal algorithm, OFU-MNL+, achieving a matching upper bound of $\mathcal{O}(\sqrt{v_{0}K}/(v_{0}+K)d\sqrt{T})$ (Theorem 2) up to logarithmic factors, while requiring only a constant computation cost per round. The results indicate that the regret improves as the assortment size $K$ increases, unless $v_{0}=\Theta(K)$. To the best of our knowledge, this is the first study to demonstrate the dependence of regret on the attraction parameter for the outside option $v_{0}$ and to highlight the advantages of a larger assortment size $K$ which aligns with intuition. That is, this is the first work to show that a regret upper bound (in either contextual or non-contextual setting) decreases as $K$ increases. Additionally, we provide instance-dependent minimax regret bounds (Proposition 1 and 2), up to logarithmic factors.
  • Under non-uniform rewards, with setting $v_{0}=\Theta(1)$ following the convention in contextual MNL bandits [5, 16, 41, 7, 38, 39, 8, 42, 4, 51, 31], we establish a regret lower bound of $\Omega(d\sqrt{T})$ (Theorem 3). To the best of our knowledge, this is the first and tightest lower bound established under non-uniform rewards. Moreover, OFU-MNL+ also achieves a matching upper bound (up to logarithmic factors) of $\tilde{\mathcal{O}}(d\sqrt{T})$ (Theorem 4) in this setting.
  • We also conduct numerical experiments and show that our algorithm consistently outperforms the existing MNL bandit algorithms while maintaining a constant computation cost per round. Furthermore, the empirical results corroborate our theoretical findings regarding the dependence of regret on the reward structure, $v_{0}$ and $K$.
Overall, our paper addresses the long-standing open problem of closing the gap between upper and lower bounds for contextual MNL bandits. Our proposed algorithm is the first to achieve both provably optimality (up to logarithmic factors) and practicality with improved computation.
2 Related Work
Lower bounds of MNL bandits. In contextual MNL bandits, to the best of our knowledge, only Chen et al. [14] proved a lower bound of $\Omega(d\sqrt{T}/K)$ with the attraction parameter for the outside option set at $v_{0}=1$. However, in the non-contextual setting, there exist improved lower bounds in terms of $K$. Agrawal et al. [7] demonstrated a lower bound of $\Omega(\sqrt{NT/K})$, and Chen and Wang [13] established a lower bound of $\Omega(\sqrt{NT})$. By setting $d=N$, one can derive equivalent lower bounds for the contextual setting, specifically $\Omega(\sqrt{dT/K})$ and $\Omega(\sqrt{dT})$, respectively. However, Agrawal et al. [7] and Chen and Wang [13] assumed $v_{0}=K$ when establishing their lower bounds, which differs from the setting used by Chen et al. [14], where $v_{0}=1$. Moreover, to the best of our knowledge, all existing works Chen et al. [14], Agrawal et al. [7], Chen and Wang [13] have established the lower bounds under uniform rewards. Consequently, it remains unclear what the optimal regret is, depending on the value of $v_{0}$ and the reward structure.
Upper bounds of contextual MNL bandits. Ou et al. [41] formulated a linear utility model and achieved $\tilde{\mathcal{O}}(dK\sqrt{T})$ regret; however, they assumed that utilities are fixed over time. Chen et al. [14] considered contextual MNL bandits with changing and stochastic contexts, establishing a regret of $\tilde{\mathcal{O}}(d\sqrt{T}+d^{2}K^{2})$. However, they encountered computational issues due to the need to enumerate all possible ($N$ choose $K$) assortments. To address this, Oh and Iyengar [39] proposed a polynomial-time assortment optimization algorithm, which maintains the confidence bounds in the parameter space and then calculates the upper confidence bounds of attraction parameter for each item, achieving a regret of $\tilde{\mathcal{O}}(d\sqrt{T}/\kappa)$, where $1/\kappa=\mathcal{O}(K^{2})$ is a problem-dependent constant. Perivier and Goyal [42] considered the adversarial context and uniform reward setting and improved the dependency on $\kappa$ to $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$. However, their algorithm is intractable. Agrawal et al. [4] considered a uniform rewards setting (with $v_{0}=1$) and achieved a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$. However, due to significant technical errors in their paper (refer Appendix L), we do not include a comparison with their results in this work.
Recently, Zhang and Sugiyama [51] utilized an online parameter update to construct a constant time algorithm. However, they consider a multiple-parameter choice model in which the learner estimates $K$ parameters and shares the contextual information $x_{t}$ across the items in the assortment. This model differs from ours; we use a single-parameter choice model with varying the context for each item in the assortment. Additionally, they make a stronger assumption regarding the reward than we do (see Assumption 1). Moreover, while they fix the assortment size at $K$, we allow it to be smaller than or equal to $K$. On the other hand, Zhang and Luo [50] considered a general function approximation, achieving a regret bound of $\tilde{\mathcal{O}}(K^{2.5}\sqrt{dNT})$. However, this bound scales with $K$ and $N$, and the proposed algorithm is not tractable. To the best of our knowledge, all existing methods fail to show that the regret upper bound can improve as the assortment size $K$ increases.
3 Problem Setting
Notations. For a positive integer, $n$, we denote $n]:=\{1,2,\ldots,n\}$. For a real-valued matrix $A$, we denote $\|A\|_{2}:=\sup_{x:\|x\|_{2}=1}\|Ax\|_{2}$ as the maximum singular value of $A$. For two symmetric matrices, $V$ and $W$ of the same dimensions, $V\succeq W$ means that VWV-W is positive semi-definite. Finally, we define $\mathcal{S}$ to be the set of candidate assortment with size constraint at most $K$, i.e., $\mathcal{S}=\{S\subseteq[N]:|S|\leq K\}$. While, for simplicity, we consider both $\mathcal{S}$ and the set of items $[N]$ to be stationary in this paper, it is important to note that both $\mathcal{S}$ and $[N]$ can vary over time.
Contextual MNL bandits. We consider a sequential assortment selection problem which is defined as follows. At each round $t$, the agent observes feature vectors $x_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Based on this contextual information, the agent presents an assortment $S_{t}=\{i_{1},\dots,i_{l}\}\in\mathcal{S}$, where $l\leq K$, and then observes the user purchase decision $c_{t}\in S_{t}\cup\{0\}$, where $\{0\}$ represents the “outside option” which indicates that the user did not select any of the items in $S_{t}$. The distribution of these selections follows a multinomial logistic (MNL) choice model [35], where the probability of choosing any item $i_{k}\in S_{t}$ (or the outside option) is defined as:
$p_{t}(i_{k}|S_{t},\mathbf{w}^{\star}):=\frac{\exp(x_{ti_{k}}^{\top}\mathbf{w}^{\star})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})},\quad p_{t}(0|S_{t},\mathbf{w}^{\star}):=\frac{v_{0}}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})},$ (1)
where $v_{0}$ is a known attraction parameter for the outside option and $\mathbf{w}^{\star}\in\mathbb{R}^{d}$ is an unknown parameter.
Remark 1.
In the existing literature on MNL bandits, it is commonly assumed that $v_{0}=1$ [38, 39, 42, 4, 51]. On the other hand, Chen and Wang [13], Agrawal et al. [7] assume that $v_{0}=K$  [^1] Chen and Wang [13] indeed set $v_{0}=1$ and $v_{1},\dots,v_{N}=\Theta(1/K)$. However, this is equivalent to the setting with $v_{0}=K$ and $v_{1},\dots,v_{N}=\Theta(1)$. to induce a tighter lower bound in terms of $K$. Later, we will explore how these differing assumptions create fundamentally different problems, leading to different regret lower bounds (Subsection 5.1).
The choice response for each item $i\in S_{t}\cup\{0\}$ is defined as $y_{ti}:=\mathbbm{1}(c_{t}=i)\in\{0,1\}$. Hence, the choice feedback variable $\mathbf{y}_{t}:=(y_{t0},y_{ti_{1}},\dots y_{ti_{l}})$ is sampled from the following multinomial (MNL) distribution: $\mathbf{y}_{t}\sim\operatorname{MNL}\{1,(p_{t}(0|S_{t},\mathbf{w}^{\star}),\dots,p_{t}(i_{l}|S_{t},\mathbf{w}^{\star}))\}$, where the parameter 1 indicates that $\mathbf{y}_{t}$ is a single-trial sample, i.e., $y_{t0}+\sum_{k=1}^{l}y_{ti_{k}}=1$. For each $i\in S_{t}\cup\{0\}$, we define the noise $\epsilon_{ti}:=y_{ti}-p_{t}(i|S_{t},\mathbf{w}^{\star})$. Since each $\epsilon_{ti}$ is a bounded random variable in $[0,1]$, $\epsilon_{ti}$ is $1/4$-sub-Gaussian. At every round $t$, the reward $r_{ti}$ for each item $i$ is also given. Then, we define the expected revenue of the assortment $S$ as
$R_{t}(S,\mathbf{w}^{\star}):=\sum_{i\in S}p_{t}(i|S,\mathbf{w}^{\star})r_{ti}=\frac{\sum_{i\in S}\exp(x_{ti}^{\top}\mathbf{w}^{\star})r_{ti}}{v_{0}\!+\!\sum_{j\in S}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}$
and define $S_{t}^{\star}$ as the offline optimal assortment at time $t$ when $\mathbf{w}^{\star}$ is known a prior, i.e., $S_{t}^{\star}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}R_{t}(S,\mathbf{w}^{\star})$. Our objective is to minimize the cumulative regret over the $T$ periods:
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\sum_{t=1}^{T}R_{t}(S_{t}^{\star},\mathbf{w}^{\star})-R_{t}(S_{t},\mathbf{w}^{\star}).$
When $K=1$, $r_{t1}=1$, and $v_{0}=1$, the MNL bandit recovers the binary logistic bandit with $R_{t}(S=\{x\},\mathbf{w}^{\star})=\sigma\left(x^{\top}\mathbf{w}^{\star}\right)=1/(1+\exp(-x^{\top}\mathbf{w}^{\star}))$, where $\sigma(\cdot)$ is the sigmoid function.
Consistent with previous works on MNL bandits [39, 42, 4, 51], we make the following assumptions:
Assumption 1 (Bounded assumption).
We assume that $\|\mathbf{w}^{\star}\|_{2}\leq 1$, and for all $t\geq 1$, $i\in[N]$, $\|x_{ti}\|_{2}\leq 1$ and $r_{ti}\in[0,1]$.
Assumption 2 (Problem-dependent constant).
There exists $\kappa>0$ such that for every item $i\in S$, any $S\in\mathcal{S}$, and all round $t$, $\min_{\mathbf{w}\in\mathcal{W}}p_{t}(i|S,\mathbf{w})p_{t}(0|S,\mathbf{w})\geq\kappa$, where $\mathcal{W}=\{\mathbf{w}\in\mathbb{R}^{d}\mid\|\mathbf{w}\|_{2}\leq 1\}$.
In Assumption 1, we assume that the reward for each item $i$ is bounded by a constant, allowing the norm of the reward vector to depend on $K$, e.g., $\|\bm{\rho}_{t}\|_{2}\leq\sqrt{K}$. In contrast, Zhang and Sugiyama [51] assume that the norm of the reward vector $\bm{\rho}_{t}=[r_{t1},\dots r_{t|S_{t}|}]^{\top}\in\mathbb{R}^{|S_{t}|}$ is bounded by a constant, independent of $K$, e.g., $\|\bm{\rho}_{t}\|_{2}\leq 1$. Thus, our assumption regarding rewards is weaker than theirs.
Assumption 2 is common in contextual MNL bandits [14, 39, 42, 51]. Note that $1/\kappa$ depends on the maximum size of the assortment $K$, i.e., $1/\kappa=\mathcal{O}(K^{2})$. One of the primary goals of this paper is to show that as the assortment size $K$ increases, we can achieve an improved (or at least not worsened) regret bound. To this end, we design a dynamic assortment policy that enjoys improved dependence on $\kappa$. Note that our algorithm does not need to know $\kappa$ a priori, whereas Oh and Iyengar [38, 39] do.
4 Existing Gap between Upper and Lower Bounds in MNL Bandits
The primary objective of this paper is to establish minimax regrets in contextual MNL bandits. To explore the optimality of regret, we analyze how it depends on the attraction parameter for the outside option $v_{0}$, the maximum assortment size $K$, and the structure of rewards.
[…]
5 Algorithms and Main Results
In this section, we begin by proving the tightest regret lower bound under uniform rewards (Subsection 5.1), explicitly showing the dependence on the attraction parameter for the outside option $v_{0}$. We then introduce OFU-MNL+, an algorithm that achieves minimax optimality, up to logarithmic factors, under uniform rewards (Subsection 5.2). Notably, OFU-MNL+ is designed for efficiency, requiring only constant computation and storage costs. Finally, we establish the tightest regret lower bound and a matching minimax optimal regret upper bound under non-uniform rewards (Subsection 5.3).
5.1 Regret Lower Bound under Uniform Rewards
In this subsection, we present a lower bound for the worst-case expected regret in the uniform reward setting ($r_{ti}=1$). This covers all applications where the objective is to maximize the appropriate “click-through rate” by offering the assortment.
Theorem 1 (Regret lower bound, Uniform rewards).
Let $d$ be divisible by $4$ and let Assumption 1 hold true. Suppose $T\geq C\cdot d^{4}(v_{0}+K)/K$ for some constant $C>0$. Then, in the uniform reward setting, for any policy $\pi$, there exists a worst-case problem instance such that the worst-case expected regret of $\pi$ is lower bounded as follows:
$\sup_{\mathbf{w}}\mathbb{E}^{\pi}_{\mathbf{w}}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]=\Omega\left(\frac{\sqrt{v_{0}K}}{v_{0}+K}\cdot d\sqrt{T}\right).$
[…]
5.2 Minimax Optimal Regret Upper Bound under Uniform Rewards
In this subsection, we propose a new algorithm OFU-MNL+, which enjoys minimax optimal regret up to logarithmic factors in the case of uniform rewards. Note that, since the revenue is an increasing function when rewards are uniform, maximizing the expected revenue $R_{t}(S,\mathbf{w})$ over all $S\in\mathcal{S}$ always yields exactly $K$ items, i.e., $|S_{t}|=|S_{t}^{\star}|=K$.
Our first step involves constructing the confidence set for the online parameter.
Online parameter estimation. Instead of performing MLE as in previous works [14, 39, 42], inspired by Zhang and Sugiyama [51], we use the online mirror descent algorithm to estimate parameter. We first define the multinomial logistic loss function at round $t$ as:
$\ell_{t}(\mathbf{w}):=-\sum_{i\in S_{t}}y_{ti}\log p_{t}(i|S_{t},\mathbf{w}).$ (2)
In Proposition C.1, we will show that the loss function has the constant parameter self-concordant-like property. We estimate the true parameter $\mathbf{w}^{\star}$ as follows:
$\mathbf{w}_{t+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\langle\nabla\ell_{t}(\mathbf{w}_{t}),\mathbf{w}\rangle+\frac{1}{2\eta}\|\mathbf{w}-\mathbf{w}_{t}\|_{\tilde{H}_{t}}^{2}\,,\quad\forall t\geq 1,$ (3)
where $\eta>0$ is the step-size parameter to be specified later, and $\tilde{H}_{t}:=H_{t}+\eta\mathcal{G}_{t}(\mathbf{w}_{t})$, where
$\mathcal{G}_{t}(\mathbf{w}):=\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w})x_{ti}x_{ti}^{\top}-\sum_{i\in S_{t}}\sum_{j\in S_{t}}p_{t}(i|S_{t},\mathbf{w})p_{t}(j|S_{t},\mathbf{w})x_{ti}x_{tj}^{\top},$
and $H_{t}:=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Note that $\mathcal{G}_{t}(\mathbf{w})=\nabla^{2}\ell_{t}(\mathbf{w})$. This online estimator is efficient in terms of both computation and storage. By a standard online mirror descent formulation [40], the optimization problem in Equation (3) can be solved using a single projected gradient step through the following equivalent formula:
$\mathbf{w}^{\prime}_{t+1}=\mathbf{w}_{t}-\eta\tilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t}),\quad\text{and}\quad\mathbf{w}_{t+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\|\mathbf{w}-\mathbf{w}^{\prime}_{t+1}\|_{\tilde{H}_{t}},$ (4)
which enjoys a computational cost of only $\mathcal{O}(Kd^{3})$, completely independent of $t$ [36, 51]. Regarding storage costs, the estimator does not need to store all historical data because both $\tilde{H}_{t}$ and $H_{t}$ can be updated incrementally, requiring only $\mathcal{O}(d^{2})$ storage.
Furthermore, the estimator allows for a $\kappa$-independent confidence set, leading to an improved regret.
Lemma 1 (Online parameter confidence set).
Let $\delta\in(0,1]$. Under Assumption 1, with $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda=84\sqrt{2}d\eta$, we define the following confidence set
$\mathcal{C}_{t}(\delta):=\{\mathbf{w}\in\mathcal{W}\mid\|\mathbf{w}_{t}-\mathbf{w}\|_{H_{t}}\leq\beta_{t}(\delta)\},$
where $\beta_{t}(\delta)=\mathcal{O}\left(\sqrt{d}\log t\log K\right)$. Then, we have $\textup{Pr}[\forall t\geq 1,\mathbf{w}^{\star}\in\mathcal{C}_{t}(\delta)]\geq 1-\delta$.
Armed with the online estimator, we construct the computationally efficient optimistic revenue.
Algorithm 1 OFU-MNL+
1:Inputs: regularization parameter $\lambda$, probability $\delta$, confidence radius $\beta_{t}(\delta)$, step size $\eta$.
2:Initialize: $H_{1}=\lambda\mathbf{I}_{d}$ and $\mathbf{w}_{1}$ as any point in $\mathcal{W}$.
3:for round $t=1,2,\cdots,T$ do
4:  Compute $\alpha_{ti}=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$ for all $i\in[N]$.
5:  Offer $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$ and observe $\mathbf{y}_{t}$.
6:  Update $\tilde{H}_{t}=H_{t}+\eta\mathcal{G}_{t}(\mathbf{w}_{t})$, and update the estimator $\mathbf{w}_{t+1}$ by Equation (3).
7:  Update $H_{t+1}=H_{t}+\mathcal{G}_{t}(\mathbf{w}_{t+1})$.
8:end for
Computationally efficient optimistic expected revenue. To balance the exploration and exploitation trade-off, we use the upper confidence bounds (UCB) technique, which have been widely studied in many bandit problems, including $K$-arm bandits [9, 30] and linear bandits [1, 18].
At each time $t$, given the confidence set in Lemma 1, we first calculate the optimistic utility $\alpha_{ti}$ as:
$\alpha_{ti}:=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}},\quad\,\forall i\in[N].$ (5)
The optimistic utility $\alpha_{ti}$ is composed of two parts: the mean utility estimate $x_{ti}^{\top}\mathbf{w}_{t}$ and the standard deviation $\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$. In the proof of the regret upper bound, we show that $\alpha_{ti}$ serves as an upper bound for $x_{ti}^{\top}\mathbf{w}^{\star}$, assuming that the true parameter $\mathbf{w}^{\star}$ falls within the confidence set $\mathcal{C}_{t}(\delta)$. Based on $\alpha_{ti}$, we construct the optimistic expected revenue for the assortment $S$, defined as follows:
$\tilde{R}_{t}(S):=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})},$ (6)
where $r_{ti}=1$. Then, we offer the set $S_{t}$ that maximizes the optimistic expected revenue, $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$. Given our assumption that all rewards are of unit value, the optimization problem is equivalent to selecting the $K$ items with the highest optimistic utility $\alpha_{ti}$. Consequently, solving the optimization problem incurs a constant computational cost of $\mathcal{O}(N)$.
Remark 2 (Comparison to Zhang and Sugiyama [51]).
In Zhang and Sugiyama [51], the MNL choice model is outlined with a shared context $x_{t}$ and distinct parameters $\mathbf{w}_{1}^{\star},\dots,\mathbf{w}_{K}^{\star}$ for each choice. Conversely, our model employs a single parameter $\mathbf{w}^{\star}$ across all choices and has varying contexts for each item in the assortment $S$, $x_{t1},\dots x_{ti_{|S|}}$. Due to this discrepancy in the choice model, directly applying Proposition 1 from Zhang and Sugiyama [51], which constructs the optimistic revenue by adding bonus terms to the estimated revenue, incurs an exponential computational cost in our problem setting. This complexity arises because the optimistic revenue must be calculated for every possible assortment $S\in\mathcal{S}$; therefore, it is necessary to enumerate all potential assortments ($N$ choose $K$) to identify the one that maximizes the optimistic revenue As a result, extending the approach in Zhang and Sugiyama [51] to our setting is non-trivial, requiring a different analysis.
We now present the regret upper bound of OFU-MNL+ in the uniform reward setting.
Theorem 2 (Regret upper bound of OFU-MNL+, Uniform rewards).
Let $\delta\in(0,1]$ and Assumptions 1 and 2 hold. In the uniform reward setting, by setting $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda=84\sqrt{2}d\eta$, with probability at least $1-\delta$, the cumulative regret of OFU-MNL+ is upper-bounded by
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(\frac{\sqrt{v_{0}K}}{v_{0}+K}\cdot d\sqrt{T}+\frac{1}{\kappa}d^{2}\right).$
Discussion of Theorem 2. If $T\geq\mathcal{O}(d^{2}(v_{0}+K)^{2}/(\kappa^{2}v_{0}K))$, Theorem 2 shows that our algorithm OFU-MNL+ achieves minimax optimal regret (up to logarithmic factor) in terms of all $d$, $T$, $K$, and even $v_{0}$. To the best of our knowledge, ignoring logarithmic factors, our proposed algorithm is the first computationally efficient, minimax optimal algorithm in (adversarial) contextual MNL bandits. When $v_{0}=\Theta(1)$, which is the convention in existing MNL bandit literature [38, 39, 42, 4, 51], OFU-MNL+ obtains $\tilde{\mathcal{O}}(d\sqrt{T/K})$ regret. This represents an improvement over the previous upper bound of Perivier and Goyal [42] 222 Perivier and Goyal [42] also consider the uniform rewards ($r_{ti}=1$) with $v_{0}=1$., which is $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$, by a factor of $K$. This improvement can be attributed to two key factors: an improved, constant, self-concordant-like property of the loss function (Proposition C.1) and a $K$-free elliptical potential lemma (Lemma E.2). Furthermore, by employing an improved bound for the second derivative of the revenue (Lemma E.3), we achieve an enhancement in the regret for the second term, $d^{2}/\kappa$, by a factor of $K^{4}$, in comparison to Perivier and Goyal [42]. Unless $v_{0}=\Theta(K)$, Theorem 2 indicates that the regret decreases as the assortment size $K$ increases. To the best of our knowledge, this is the first algorithm in MNL bandits to show that increasing $K$ results in a reduction in regret. Moreover, when reduced to the logistic bandit, i.e., $K=1$, $r_{t1}=1$, and $v_{0}=1$, our algorithm can also achieve a regret of $\tilde{\mathcal{O}}(d\sqrt{\kappa T})$ by Corollary 1 in Zhang and Sugiyama [51], which is consistent with the results in Abeille et al. [3], Faury et al. [21]. The proof is deferred to Appendix E.
Remark 3 (Efficiency of OFU-MNL+).
The proposed algorithm is computationally efficient in both parameter updates and assortment selections. Since we employ online parameter estimation, akin to Zhang and Sugiyama [51], our algorithm demonstrates computational efficiency in parameter estimation, incurring only incurring $\mathcal{O}(Kd^{3})$ computation cost and $\mathcal{O}(d^{2})$ storage cost, which are completely independent of $t$. Furthermore, a naive approach to selecting the optimistic assortment requires enumerating all possible ($N$ choose $K$) assortments, resulting in exponential computational cost [14]. However, by constructing the optimistic expected revenue according to Equation (6) (inspired by Oh and Iyengar [39]), our algorithm needs only $\mathcal{O}(N)$ computational cost.
5.3 Regret Upper & Lower Bounds under Non-Uniform Rewards
In this subsection, we propose regret upper and lower bounds in the non-uniform reward setting. In this scenario, the sizes of both the chosen assortment $S_{t}$, and the optimal assortment, $S_{t}^{\star}$ are not fixed at $K$. Therefore, we cannot guarantee an improvement in regret even as $K$ increases.
We first prove the regret lower bound in the non-uniform reward setting.
Theorem 3 (Regret lower bound, Non-uniform rewards).
Suppose Assumption 1 holds and $T\geq C\cdot d^{4}(v_{0}+1)/1$ for some constant $C>0$. Let $d$ be divisible by $4$ and $v_{0}=\Theta(1)$. Then, under the non-uniform reward setting, for any policy $\pi$, there exists an instance such that the worst-case expected regret of $\pi$ is lower bounded as follows:
$\sup_{\mathbf{w}}\mathbb{E}^{\pi}_{\mathbf{w}}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]=\Omega\left(d\sqrt{T}\right).$
Discussion of Theorem 3. In contrast to Theorem 1, which considers uniform rewards, the regret lower bound is independent of the assortment size $K$. Note that Theorem 3 does not claim that non-uniform rewards inherently make the problem more difficult. Rather, it implies that there exists an instance with adversarial non-uniform rewards, where regret does not improve even with an increase in $K$. Moreover, the assumption that $v_{0}=\Theta(1)$ is common in the existing literature on contextual MNL bandits [38, 39, 42, 4, 51] (refer Appendix C.1). To the best of our knowledge, this is the first established lower bound for non-uniform rewards in MNL bandits, even in the non-contextual setting. The proof is deferred to Appendix G.
We also prove a matching upper bound up to logarithmic factors. The algorithm OFU-MNL+ is also applicable in the case of non-uniform rewards. However, because the optimistic expected revenue $\tilde{R}_{t}(S)$ is no longer an increasing function of $\alpha_{ti}$, optimizing for $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$ no longer equates to simply selecting the top $K$ items with the highest optimistic utility. Instead, we employ assortment optimization methods introduced in Rusmevichientong et al. [45], Davis et al. [19], which are efficient polynomial-time (independent of $t$)  [^3] An interior point method would generally solve the problem with a computational complexity of $\mathcal{O}(N^{3.5})$. algorithms available for solving this optimization problem. Therefore, our algorithm is also computationally efficient under non-uniform rewards.
Theorem 4 (Regret upper bound of OFU-MNL+, Non-uniform rewards).
Under the same assumptions and parameter settings as Theorem 2, if the rewards are non-uniform and $v_{0}=\Theta(1)$, then with a probability at least $1-\delta$, the cumulative regret of OFU-MNL+ is upper-bounded by
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})$ $=\tilde{\mathcal{O}}\left(d\sqrt{T}+\frac{1}{\kappa}d^{2}\right).$
Discussion of Theorem 4. If $T\geq\mathcal{O}(d^{2}/\kappa^{2})$, our algorithm achieves a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$ when the reward for each item is non-uniform, demonstrating that OFU-MNL+ is minimax optimal up to a logarithmic factor. Recall that we relax the bounded assumption on the reward compared to Zhang and Sugiyama [51] (refer Assumption 1); thus, we allow the sum of the squared rewards in the assortment to scale with $K$. Consequently, we need a novel approach to achieve the regret that does not scale with $K$. To this end, we centralize the features, i.e., $\tilde{x}_{ti}=x_{ti}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}_{t+1})}[x_{tj}]$, and propose a novel elliptical potential lemma for them, as detailed in Lemma H.3. Note that our algorithm is capable of achieving $1/\kappa$-free regret (in the leading term) under both uniform and non-uniform rewards. In contrast, the algorithm in Perivier and Goyal [42] is limited to achieving this only in the uniform reward setting. Furthermore, compared to the regret bound in Chen et al. [14], which is $\tilde{\mathcal{O}}(d\sqrt{T})$, our regret bounds has the same order of regret with theirs. However, their algorithm is computationally intractable as it requires enumerating all possible assortments, whereas our algorithm incurs only a constant computational cost per round. The proof is deferred to Appendix H.
6 Instance-Dependent Bounds
In this section, we show that instance-dependent upper and lower bounds are also achievable under uniform rewards. We define the degree of non-linearity for the optimal assortment $S^{\star}_{t}$ at round $t$ under the true parameter $\mathbf{w}^{\star}$ as $\kappa^{\star}_{t}:=\sum_{i\in S^{\star}_{t}}p_{t}(i|S^{\star}_{t},\mathbf{w}^{\star})p_{t}(0|S^{\star}_{t},\mathbf{w}^{\star})$. We first establish the instance-dependent lower bound under uniform rewards.
Proposition 1 (Instance-dependent regret lower bound, Uniform rewards).
Under the same conditions as Theorem 1, for any policy $\pi$ and for $T\geq d^{2}/\kappa$, there exists a worst-case problem instance such that the worst-case expected regret of $\pi$ is lower bounded as follows:
$\sup_{\mathbf{w}}\mathbb{E}^{\pi}_{\mathbf{w}}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]=\Omega\left(d\sqrt{\sum_{t=1}^{T}\kappa^{\star}_{t}}\right).$
The proof is deferred to Appendix I. We also provide the matching upper bound.
Proposition 2 (Instance-dependent regret upper bound of OFU-MNL+, Uniform rewards).
Under the same assumptions, parameter settings, and reward structure as Theorem 2, with a probability at least $1-\delta$, the cumulative regret of OFU-MNL+ is upper-bounded by
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})$ $=\tilde{\mathcal{O}}\left(d\sqrt{\sum_{t=1}^{T}\kappa^{\star}_{t}}+\frac{1}{\kappa}d^{2}\right).$
The proof is deferred to Appendix J. For sufficiently large $T$, the regret upper bound (Proposition 2) matches the regret lower bound (Proposition 1), up to logarithmic factor. To the best of our knowledge, these are the first minimax instance-dependent regret bounds under uniform rewards. Note that, in the worst case, $\kappa^{\star}_{t}=\tilde{\mathcal{O}}(\sqrt{v_{0}K}/(v_{0}+K))$, which indicates that these results provide a strict improvement over the worst-case bounds given in Theorems 1 and 2.
Some readers may expect instance-dependent regret bounds for non-uniform rewards as well. Unfortunately, we were unable to establish these. Recall that $\kappa^{\star}_{t}$ represents the degree of non-linearity for the optimal assortment $S^{\star}_{t}$. However, in the proofs for regret bounds, we encounter terms associated with the chosen assortment $S_{t}$, such as $\sum_{i\in S^{\star}_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})$. To address this, we use the mean value theorem-based analysis (Lemma I.5) to replace this quantity with $\kappa^{\star}_{t}$. Under non-uniform rewards, however, the mean value theorem does not apply because the sizes and rewards of $S^{\star}_{t}$ and $S_{t}$ may differ. Addressing this problem would be an interesting direction for future research.
7 Numerical Experiments
Figure 1: Cumulative regret (left three, $K=5,10,15$) and runtime per round (rightmost one, $K=15$) under uniform rewards (first row) and non-uniform rewards (second row) with $v_{0}=1$.
In this section, we empirically evaluate the performance of our algorithm OFU-MNL+. We measure cumulative regret over $T=3000$ rounds. For each experimental setup, we run the algorithms across $20$ independent instances and report the average performance. In each instance, the underlying parameter $\mathbf{w}^{\star}$ is randomly sampled from a $d$-dimensional uniform distribution, where each element of $\mathbf{w}^{\star}$ lies within the range $1/\sqrt{d},1/\sqrt{d}]$ and is not known to the algorithms. Additionally, the context features $x_{ti}$ are drawn from a $d$-dimensional multivariate Gaussian distribution, with each element of $x_{ti}$ clipped to the range $1/\sqrt{d},1/\sqrt{d}]$. This setup ensures compliance with Assumption 1. In the uniform reward setting (first row of Figure 1), the combinatorial optimization step to choose the assortment reduces to sorting items by their utility estimate. In the non-uniform reward setting (second row of Figure 1), the rewards are sampled from a uniform distribution in each round, i.e., $r_{ti}\sim\operatorname{Unif}(0,1)$. Refer Appendix K for more details.
We compare the performance of OFU-MNL+ with those of the practical and state-of-the-art algorithms: the Upper Confidence Bound-based algorithm, UCB-MNL [38], and the Thompson Sampling-based algorithm, TS-MNL [38]. Figure 1 demonstrates that our algorithm significantly outperforms other baseline algorithms. In the uniform reward setting, as $K$ increases, the cumulative regrets of all algorithms tend to decrease. In contrast, this trend is not observed in the non-uniform reward setting. Furthermore, the results also show that our algorithm maintains a constant computation cost per round, while the other algorithms exhibit a linear dependence on $t$. In Appendix K, we present the additional runtime curves (Figure K.1) as well as the regret curves of the other configuration where $v_{0}=\Theta(K)$ (Figure K.2). All of these empirical results align with our theoretical results.
8 Conclusion
[…]
Acknowledgements
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. 2022R1C1C1006859, 2022R1A4A1030579, and RS-2023-00222663) and by AI-Bio Research Grant through Seoul National University.
Appendix A Further Related Work
In this section, we discuss additional related works that complement Section 2. For simplicity, we consider only the dependence on the number of rounds $t$ for a computation cost in big-$\mathcal{O}$ notation.
[…]
Multinomial Logistic (MNL) Bandits. There are two main approaches to multinomial logistic (MNL) bandits: the multiple-parameter choice model and the single-parameter choice model. In the multiple-parameter choice model, the learner estimates parameters for each choice in the assortment ($\mathbf{w}_{1}^{\star},\dots,\mathbf{w}_{K}^{\star}$) with a shared context $x_{t}$. In this setting, Amani and Thrampoulidis [8] proposed a feasible algorithm that achieves a regret upper bound of $\tilde{\mathcal{O}}(dK\sqrt{\kappa T})$ with a $\mathcal{O}(t)$ computation cost. They also proposed an intractable algorithm that achieves an improved regret of $\tilde{\mathcal{O}}(dK^{3/2}\sqrt{T})$. Zhang and Sugiyama [51] introduced a computationally and statistically efficient algorithm that obtains a regret of $\tilde{\mathcal{O}}(dK\sqrt{T})$. Recently, Lee et al. [31] further improved the regret by a factor of $\sqrt{K}$, achieving $\tilde{\mathcal{O}}(d\sqrt{KT})$ regret. In the multiple-parameter case, the regret’s dependence on $K$ is unavoidable since the number of unknown parameters depends on $K$.
On the other hand, the single-parameter choice model, closely related to ours, shares the parameter $\mathbf{w}^{\star}$ cross the choices, with varying contexts for each choice. The learner offers a set of items $S_{t}$, with $|S_{t}|\leq K$ at each round. This setting involves a combinatorial optimization to choose the assortment $S_{t}$, making it more challenging to devise a tractable algorithm. As extensively discussed in Section 2, no previous studies have definitively confirmed whether the existing lower or upper bounds are tight. As shown in Table 1, many studies have presented their results in inconsistent settings with varying reward structures and values of $v_{0}$, adding to the ambiguity about the bounds’ optimality. In this paper, we address these issues by bridging the gap between the lower and upper bounds of regret through a careful categorization of the settings. We propose an algorithm that is both provably optimal, up to logarithmic factors, and computationally efficient, significantly enhancing the theoretical and practical understanding of MNL bandits.
Generalized Linear Bandits. In generalized linear bandits [22, 25, 33, 2, 29, 26, 27, 32], the expected rewards are modeled using a generalized linear model. These problems generalize logistic bandits by incorporating a general exponential family link function instead of the logistic link function. The algorithms proposed for generalized linear bandits also exhibit a dependence on the nonlinear term $\kappa$. However, our problem setting (single-parameter MNL bandits) considers a more complex state space where multiple arms are pulled simultaneously.
Combinatorial Bandits. Another related stream of literature is combinatorial bandits [12, 43, 28, 52, 44, 34], particularly top-$k$ combinatorial bandits [44]. In top-$k$ combinatorial bandits, the decision set includes all subsets of size $k$ out of $n$ arms, and the reward for each action is the sum of the rewards of the $k$ selected arms. In this framework, the rewards are assumed to be independent of the entire set of arms played in round $t$. In contrast, in our setting, the reward for each individual arm depends on the whole set of arms played.
Recently, Choi et al. [17], Cao and Sun [11] have considered the cascading assortment bandit problem, which encompasses the MNL bandit problem as a special case where the cascading length is $1$. However, these works do not strictly encompass our results. Choi et al. [17] only consider uniform rewards, and achieve a regret upper bound of $\tilde{\mathcal{O}}(d\sqrt{T})$, which avoids dependence on both the cascading length and $\kappa$. When the cascading length is $1$, our result (Theorem 2) improves upon theirs by a factor of $\sqrt{K}$. Moreover, their computation cost per round is $\mathcal{O}(t)$ since they employ MLE to estimate the parameter. Cao and Sun [11] consider non-uniform rewards, and achieve a regret upper bound of $\tilde{\mathcal{O}}(h^{2}d\sqrt{MT})$, where $M$ is the cascading length and $h\geq\frac{p(i|S,\mathbf{w})}{p(i|S,\mathbf{w^{\prime}})}$ for all $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{W}$, $S\in\mathcal{S}$, and $i\in S\cup\{0\}$. However, their regret bound still suffers from a harmful dependence on $h^{2}$, which can be exponentially large.
Appendix B Notation
We denote $T$ as the total number of rounds and $t\in[T]$ as the current round. We denote $N$ as the total number of items, $K$ as the maximum size of assortments, and $d$ as the dimension of feature vectors.
For notational simplicity, we define the loss function in two different forms throughout the proof:
$\ell_{t}(\mathbf{w})$ $=-\sum_{i\in S_{t}}y_{ti}\log p_{t}(i|S_{t},\mathbf{w})=-\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(x_{ti}^{\top}\mathbf{w})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w})}\right),$
$\ell(\mathbf{z}_{t},\mathbf{y}_{t})$ $=-\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(z_{ti})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(z_{tj})}\right),$
where $z_{ti}=x_{ti}^{\top}\mathbf{w}$, $\mathbf{z}_{t}=(z_{ti})_{i\in S_{t}}\in\mathbb{R}^{|S_{t}|}$, and $\mathbf{y}_{t}=(y_{ti})_{i\in S_{t}}\in\mathbb{R}^{|S_{t}|}$. Thus, $\ell_{t}(\mathbf{w})=\ell(\mathbf{z}_{t},\mathbf{y}_{t})$.
We offer a Table B.1 for convenient reference.
Table B.1: Symbols
$x_{ti}$ feature vector for item $i$ given at round $t$
$r_{ti}$ reward for item $i$ given at round $t$
$S_{t}$ assortment chosen by an algorithm at round $t$
0 outside option
$y_{ti}$ choice response for each item $i\in S_{t}\cup\{0\}$ at round $t$
$R_{t}(S,\mathbf{w}^{\star})$ $\sum_{i\in S}p_{t}(i|S,\mathbf{w}^{\star})r_{ti}$, expected revenue of the assortment $S$ at round $t$
$\ell_{t}(\mathbf{w})$ $\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(x_{ti}^{\top}\mathbf{w})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w})}\right)$, loss function at round $t$
$\ell(\mathbf{z}_{t},\mathbf{y}_{t})$ $\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(z_{ti})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(z_{tj})}\right)$, loss function at round $t$, $z_{ti}=x_{ti}^{\top}\mathbf{w}$
$\lambda$ regularization parameter
$\mathcal{G}_{t}(\mathbf{w})$ $\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w})x_{ti}x_{ti}^{\top}-\sum_{i\in S_{t}}\sum_{j\in S_{t}}p_{t}(i|S_{t},\mathbf{w})p_{t}(j|S_{t},\mathbf{w})x_{ti}x_{tj}^{\top}$
$H_{t}$ $\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$
$\tilde{H}_{t}$ $H_{t}+\eta\mathcal{G}_{t}(\mathbf{w}_{t})$
$\alpha_{ti}$ $x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$, optimistic utility for item $i$ at round $t$
$\beta_{t}(\delta)$ $\mathcal{O}\left(\sqrt{d}\log t\log K\right)$, confidence radius at round $t$
$\tilde{R}_{t}(S)$ $\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$, optimistic expected revenue for the assortment $S$ at round $t$
Appendix C Properties of MNL function
In this section, we present key properties of the MNL function and its associated loss, which are used throughout the paper.
C.1 Attraction parameter for Outside Option: $v_{0}=\Theta(1)$ is Common in Contextual MNL Bandits
In this subsection, we explain why the assumption that $v_{0}=\Theta(1)$ is made without loss of generality. Let the original feature vectors be $x^{\prime}_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Suppose that a context for the outside option $x_{t0}^{\prime}$ is given and the probability of choosing any item $i\in S_{t}\cup\{0\}$ is defined as
$p_{t}(i|S_{t},\mathbf{w}^{\star})=\frac{\exp((x^{\prime}_{ti})^{\top}\mathbf{w}^{\star})}{\sum_{j\in S_{t}\cup\{0\}}\exp((x^{\prime}_{tj})^{\top}\mathbf{w}^{\star})}.$
Then, by dividing the denominator and numerator by $\exp\left((x^{\prime}_{t0})^{\top}\mathbf{w}^{\star}\right)$, and defining $x_{ti}:=x^{\prime}_{ti}-x^{\prime}_{t0}$, we obtain the MNL probability in the form presented in Equation (1) with $v_{0}=\exp(0)=1$. Note that this division does not change the probability. Therefore, $v_{0}=\Theta(1)$ is natural and common in contextual MNL bandit literature.
C.2 Self-concordant-like Function
Definition C.1 (Self-concordant-like function, Tran-Dinh et al. 47).
A convex function $f\in\mathcal{C}^{3}(\mathbb{R}^{m})$ is $M$-self-concordant-like function with constant $M$ if:
$|\phi^{\prime\prime\prime}(s)|\leq M\|\mathbf{b}\|_{2}\phi^{\prime\prime}(s).$
for $s\in\mathbb{R}$ and $M>0$, where $\phi(s):=f(\mathbf{a}+s\mathbf{b})$ for any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{m}$.
Then, the MNL loss defined in Equation (2) is $3\sqrt{2}$-self-concordant-like function.
Proposition C.1.
For any $t\in[T]$, the multinomial logistic loss $\ell_{t}(\mathbf{w})$, defined in Equation (2), is $3\sqrt{2}$-self-concordant-like.
Proof.
Consider the function $\phi(s):=\log\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)$, where $\mathbf{a}=[a_{0},\dots,a_{n}]^{\top}\in\mathbb{R}^{n+1}$ and $\mathbf{b}=[b_{0},\dots,b_{n}]^{\top}\in\mathbb{R}^{n+1}$. Then, by simple calculus, we have
$\phi^{\prime\prime}(s)=\frac{\sum_{i<j}(a_{i}-a_{j})^{2}e^{a_{i}s+b_{i}}e^{a_{j}s+b_{j}}}{\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)^{2}}\geq 0,$
and
$\phi^{\prime\prime\prime}(s)=\frac{\sum_{i<j}(a_{i}-a_{j})^{2}e^{a_{i}s+b_{i}}e^{a_{j}s+b_{j}}\left[\sum_{k=0}^{n}(a_{i}+a_{j}-2a_{k})e^{a_{k}s+b_{k}}\right]}{\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)^{3}}.$ (C.1)
Note that for all $i,j,k=0,\dots,n$,
$|a_{i}+a_{j}-2a_{k}|\leq\sqrt{6}\sqrt{a_{i}^{2}+a_{j}^{2}+a_{k}^{2}}\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|.$ (C.2)
Therefore, we have
$\left|\sum_{k=0}^{n}(a_{i}+a_{j}-2a_{k})e^{a_{k}s+b_{k}}\right|\leq\sum_{k=0}^{n}\left|a_{i}+a_{j}-2a_{k}\right|e^{a_{k}s+b_{k}}\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|\sum_{k=0}^{n}e^{a_{k}s+b_{k}}.$ (C.3)
Plugging in Equation (C.3) into Equation (C.1), we obtain
$\phi^{\prime\prime\prime}(s)\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|\phi^{\prime\prime}(s).$ (C.4)
Now, we are ready to prove the proposition. For any $t\in[T]$, let $n=|S_{t}|$ and $c_{1}=x_{ti_{1}},c_{2}=x_{ti_{2}},\dots,c_{n}=x_{ti_{n}}$. Define a function $f\in\mathcal{C}^{3}:\mathbb{R}^{d}\rightarrow\ \mathbb{R}$ as $f(\bm{\theta}):=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}}\right)$. Let $\bm{\delta}\in\mathbb{R}^{d}$ and let $f(\bm{\theta}+s\bm{\delta})=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}+sc_{i}^{\top}\bm{\delta}}\right)=\log\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)=\phi(s)$, where $a_{i}=c_{i}^{\top}\bm{\delta}$, $b_{i}=c_{i}^{\top}\bm{\theta}$ for $i=1,\dots,n$, and $a_{i}=0$ and $b_{i}=\log v_{0}$ for $i=0$. Then, by Equation (C.4), we get
$|\phi^{\prime\prime\prime}(s)|$ $\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|\phi^{\prime\prime}(s)=3\sqrt{2}\max_{i=1,\dots,n}|c_{i}^{\top}\bm{\delta}|\phi^{\prime\prime}(s)$
$\leq 3\sqrt{2}\max_{i=1,\dots,n}\|c_{i}\|_{2}\|\bm{\delta}\|_{2}\phi^{\prime\prime}(s)\leq 3\sqrt{2}\|\bm{\delta}\|_{2}\phi^{\prime\prime}(s),$
where the last inequality holds due to Assumption 1 that $\|c_{i}\|_{2}=\|x_{tj_{i}}\|_{2}\leq 1$. Then, by Definition C.1, $f$ is $3\sqrt{2}$-self-concordant-like. Since $\ell_{t}$ is the sum of $f$ and a linear operator, which has third derivatives equal to zero, it follows that $\ell_{t}$ is also $3\sqrt{2}$-self-concordant-like function. ∎
Remark C.1.
[…]
Lemma C.1 (Theorem 3 of Tran-Dinh et al. 47).
A convex function $\ell\in\mathcal{C}^{3}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $M$-self-concordant-like if and only if for any $\mathbf{v},\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\in\mathbb{R}^{d}$, we have
$|\langle D^{3}\ell(\mathbf{v})[\mathbf{u}_{1}]\mathbf{u}_{2},\mathbf{u}_{3}\rangle|\leq M\|\mathbf{u}_{1}\|_{2}\|\mathbf{u}_{2}\|_{\nabla^{2}\ell(\mathbf{v})}\|\mathbf{u}_{3}\|_{\nabla^{2}\ell(\mathbf{v})}.$
Appendix D Proof of Theorem 1
In this section, we provide the proof of Theorem 1. The proof structure is similar to the one presented in Chen et al. [14]. However, unlike their approach, we explicitly derive a bound that includes $v_{0}$. Furthermore, by establishing a tighter upper bound for the KL divergence (Lemma D.2), we derive a bound that is tighter than the one provided by Chen et al. [14].
D.1 Adversarial Construction and Bayes Risk
Let $\epsilon\in(0,1/d\sqrt{d})$ be a small positive parameter to be specified later. For every subset $V\subseteq[d]$, we define the corresponding parameter $\mathbf{w}_{V}\in\mathbb{R}^{d}$ as $\mathbf{w}_{V}]_{j}=\epsilon$ for all $j\in V$, and $\mathbf{w}_{V}]_{j}=0$ for all $j\notin V$. Then, we consider the following parameter set
$\mathbf{w}\in\mathcal{W}:=\{\mathbf{w}_{V}:V\in\mathcal{V}_{d/4}\}:=\{\mathbf{w}_{V}:V\subseteq[d],|V|=d/4\},$
where $\mathcal{V}_{k}$ denotes the class of all subsets of $[d]$ whose size is $k$. Moreover, note that $d/4$ is a positive integer, as $d$ is divisible by $4$ by construction.
The context vectors $x_{ti}$ are constructed to be invariant across rounds $t$. For each $t$ and $U\in\mathcal{V}_{d/4}$, $K$ identical context vectors  [^4] Recall that $K$ is the maximum allowed assortment capacity. $x_{U}$ are constructed as follows:
$[x_{U}]_{j}=1/\sqrt{d}\quad\text{for}\,\,j\in U;\quad[x_{U}]_{j}=0\quad\text{for}\,\,j\notin U.$
For all $V,U\in\mathcal{V}_{d/4}$, it can be verified that $\mathbf{w}_{V}$ and $x_{U}$ satisfy the requirements of a bounded assumption (Assumption 1) as follows:
$\|\mathbf{w}_{V}\|_{2}\leq\sqrt{d\epsilon^{2}}\leq 1,\quad\|x_{U}\|_{2}\leq\sqrt{d\cdot 1/d}=1.$
Therefore, the worst-case expected regret of any policy $\pi$ can be lower bounded by the worst-case expected regret of parameters belonging to $\mathcal{W}$, which can be further lower bounded by the “average” regret over a uniform prior over $\mathcal{W}$ as follows:
$\sup_{\mathbf{w}}\mathbb{E}_{\mathbf{w}}^{\pi}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]$ $=\sup_{\mathbf{w}}\mathbb{E}_{\mathbf{w}}^{\pi}\sum_{t=1}^{T}R(S^{\star},\mathbf{w})-R(S_{t},\mathbf{w})$
$\geq\max_{\mathbf{w}_{V}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}R(S^{\star},\mathbf{w}_{V})-R(S_{t},\mathbf{w}_{V})$
$\geq\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}R(S^{\star},\mathbf{w}_{V})-R(S_{t},\mathbf{w}_{V})$
$=\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})\right].$ (D.1)
This reduces the task of lower bounding the worst-case regret of any policy to the task of lower bounding the Bayes risk of the constructed parameter set.
D.2 Main Proof of Theorem 1
Proof of Theorem 1.
For any sequence of assortments $\{S_{t}\}_{t=1}^{T}$ produced by policy $\pi$, we denote an alternative sequence $\{\tilde{S}_{t}\}_{t=1}^{T}$ that provably enjoys less regret under parameterization $\mathbf{w}_{V}$.
Let $x_{U_{1}},\dots,x_{U_{L}}$ be the distinct feature vectors contained in assortments $S_{t}$ (if $S_{t}=\emptyset$, then one may choose an arbitrary feature $x_{U}$) with $U_{1},\dots,U_{L}\in\mathcal{V}_{d/4}$. Let $U^{\star}$ be the subset among $U_{1},\dots,U_{L}$ that maximizes $x_{U}^{\top}\mathbf{w}_{V}$, i.e., $U^{\star}\in\mathop{\mathrm{argmax}}_{U\in\{U_{1},\dots,U_{L}\}}x_{U}^{\top}\mathbf{w}_{V}$, where $\mathbf{w}_{V}$ is the underlying parameter. Then, we define $\tilde{S}_{t}$ as the assortment consisting of all $K$ items corresponding to feature $x_{U^{\star}}$, i.e., $\tilde{S}_{t}=\{\underbrace{x_{U^{\star}},\dots,x_{U^{\star}}}_{K}\}$.
Since the expected revenue is an increasing function, we have the following observation:
Proposition D.1 (Proposition 1 in Chen et al. 14).
$\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})\leq\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V}).$
Proposition D.1 implies that $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})\geq\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$. Hence, it is sufficient to bound $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ instead of $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})$.
To simplify notation, we denote $\tilde{U}_{t}$ as the unique $U^{\star}\in\mathcal{V}_{d/4}$ in $\tilde{S}_{t}$. We also use $\mathbb{E}_{V}$ and $\mathbb{P}_{V}$ to denote the expected value and probability, respectively, as governed by the law parameterized by $\mathbf{w}_{V}$ and under policy $\pi$. Then, we can establish a lower bound for $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ as follows:
Lemma D.1.
Suppose $\epsilon\in(0,1/d\sqrt{d})$ and define $\delta:=d/4-|\tilde{U}_{t}\cap V|$. Then, we have
$\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\delta\epsilon}{2\sqrt{d}}.$
For any $j\in V$, define random variables $\tilde{M}_{j}:=\sum_{t=1}^{T}\mathbf{1}\{j\in\tilde{U}_{t}\}$. Then, by Lemma D.1, for all $V\in\mathcal{V}_{d/4}$, we have
$\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (D.2)
Furthermore, we define $\mathcal{V}_{d/4}^{(j)}:=\{V\in\mathcal{V}_{d/4}:j\in V\}$ and $\mathcal{V}_{d/4-1}:=\{V\subseteq[d]:|V|=d/4-1\}$. By taking the average of both sides of Equation (D.2) with respect to all $V\in\mathcal{V}_{d/4}$, we obtain
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\cdot\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right)$
$=\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{j=1}^{d}\sum_{V\in\mathcal{V}_{d/4}^{(j)}}\mathbb{E}_{V}[\tilde{M}_{j}]\right)$
$=\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4-1}}\sum_{j\notin V}\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]\right)$
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\frac{\left|\mathcal{V}_{d/4-1}\right|}{\left|\mathcal{V}_{d/4}\right|}\max_{V\in\mathcal{V}_{d/4-1}}\sum_{j\notin V}\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]\right)$
$=\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\frac{\left|\mathcal{V}_{d/4-1}\right|}{\left|\mathcal{V}_{d/4}\right|}\max_{V\in\mathcal{V}_{d/4-1}}\sum_{j\notin V}\mathbb{E}_{V}[\tilde{M}_{j}]+\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right).$
For any fixed $V$, we get $\sum_{j\notin V}\mathbb{E}_{V}[\tilde{M}_{j}]\leq\sum_{j=1}^{d}\mathbb{E}_{V}[\tilde{M}_{j}]\leq dT/4$. Also, we have $\frac{\left|\mathcal{V}_{d/4-1}\right|}{\left|\mathcal{V}_{d/4}\right|}=\binom{d}{d/4-1}/\binom{d}{d/4}=\frac{d/4}{3d/4+1}\leq\frac{1}{3}$. Consequently, we derive that
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}$ $\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{6}-\max_{V\in\mathcal{V}_{d/4-1}}\sum_{j\notin V}\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|\right).$ (D.3)
Now we bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in Equation (D.3) for any $V\in\mathcal{V}_{d/4-1}$. For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$ denote the laws under $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup j}$, respectively. Then, we have
$\left|\mathbb{E}_{P}[\tilde{M}_{j}]-\mathbb{E}_{Q}[\tilde{M}_{j}\right]|$ $\leq\sum_{t=0}^{T}t\cdot\left|P[\tilde{M}_{j}=t]-Q[\tilde{M}_{j}=t]\right|$
$\leq T\cdot\sum_{t=0}^{T}\left|P[\tilde{M}_{j}=t]-Q[\tilde{M}_{j}=t]\right|$
$\leq T\cdot\|P-Q\|_{\operatorname{TV}}\leq T\cdot\sqrt{\frac{1}{2}\operatorname{KL}(P\|Q)},$ (D.4)
where $\|P-Q\|_{\operatorname{TV}}=\sup_{A}|P(A)-Q(A)|$ | is the total variation distance between $P$ and $Q$, $\operatorname{KL}(P\|Q)=\int(\log\mathrm{d}P/\mathrm{d}Q)\mathrm{d}P$ is s the Kullback-Leibler (KL) divergence between $P$ and $Q$, and the last inequality holds by Pinsker’s inequality. Now, we bound the KL divergence term using the following Lemma.
Lemma D.2.
For any $V\in\mathcal{V}_{d/4-1}$ and $j\in[d]$, there exists a positive constant $C_{\operatorname{KL}}>0$ such that
$\operatorname{KL}(P_{V}\|Q_{V\cup\{j\}})\leq C_{\operatorname{KL}}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}\cdot\frac{\mathbb{E}_{V}[\tilde{M}_{j}]\epsilon^{2}}{d}.$
Therefore, combining Equation (D.3), Equation (D.4), and Lemma D.2, we have
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}$ $\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{6}-T\sum_{j=1}^{d}\sqrt{C_{\operatorname{KL}}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}\cdot\frac{\mathbb{E}_{V}[\tilde{M}_{j}]\epsilon^{2}}{d}}\right)$
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{6}-T\sqrt{d}\cdot\sqrt{\sum_{j=1}^{d}C_{\operatorname{KL}}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}\cdot\frac{\mathbb{E}_{V}[\tilde{M}_{j}]\epsilon^{2}}{d}}\right)$
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{6}-T\sqrt{d}\cdot\sqrt{\frac{C_{\operatorname{KL}}}{4}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}\cdot T\epsilon^{2}}\right),$
where the second inequality is due to the Cauchy-Schwartz inequality and the last inequality holds because $\sum_{j=1}^{d}\mathbb{E}_{V}[\tilde{M}_{j}]\leq dT/4$. Let $C_{\operatorname{KL}}^{\prime}=C_{\operatorname{KL}}/4$.
By setting $\epsilon=\sqrt{\frac{d}{144C_{\operatorname{KL}}^{\prime}T}\cdot\frac{(v_{0}+K)^{2}}{v_{0}K}}$, we have
$\sup_{\mathbf{w}}\mathbb{E}_{\mathbf{w}}^{\pi}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]$ $\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{6}-\sqrt{C_{\operatorname{KL}}^{\prime}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}dT\epsilon^{2}}\right)$
$=\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\sqrt{\frac{(v_{0}+K)^{2}}{v_{0}K}}\cdot\frac{1}{288\sqrt{C_{\operatorname{KL}}^{\prime}}}d\sqrt{T}$
$=\Omega\left(\frac{\sqrt{v_{0}K}}{v_{0}+K}\cdot d\sqrt{T}\right).$
This concludes the proof of Theorem 1. ∎
D.3 Proofs of Lemmas for Theorem 1
D.3.1 Proof of Lemma D.1
Proof of Lemma D.1.
Let $x=x_{V}$ and $\hat{x}=x_{\tilde{U}_{t}}$ be the corresponding context vectors. Then, we have
$\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ $=\frac{K\exp\left(x^{\top}\mathbf{w}_{V}\right)}{v_{0}+K\exp\left(x^{\top}\mathbf{w}_{V}\right)}-\frac{K\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)}{v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)}$
$=\frac{v_{0}K\left(\exp\left(x^{\top}\mathbf{w}_{V}\right)-\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}{\left(v_{0}+K\exp\left(x^{\top}\mathbf{w}_{V}\right)\right)\left(v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}$
$\geq\frac{v_{0}K\left(\exp\left(x^{\top}\mathbf{w}_{V}\right)-\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}{(v_{0}+Ke)^{2}},$ (D.5)
where the inequality holds since $\max\left\{\exp\left(x^{\top}\mathbf{w}_{V}\right),\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right\}\leq e$. To further bound the right-hand side of Equation (D.5), we use the fact that $1+a\leq e^{a}\leq 1+a+a^{2}/2$ for all $a\in[0,1]$, which can be easily shown by Taylor expansion. Thus, we get
$\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ $\geq\frac{v_{0}K\left((x-\hat{x})^{\top}\mathbf{w}_{V}-(\hat{x}^{\top}\mathbf{w}_{V})^{2}/2\right)}{(v_{0}+Ke)^{2}}$
$\geq\frac{v_{0}K\left(\delta\epsilon/\sqrt{d}-(\sqrt{d}\epsilon)^{2}/2\right)}{(v_{0}+Ke)^{2}}$
$\geq\frac{v_{0}K\delta\epsilon}{2\sqrt{d}(v_{0}+Ke)^{2}},$
where the last inequality holds because $(\sqrt{d}\epsilon)^{2}\leq\delta\epsilon/\sqrt{d}$ when $\epsilon\in(0,1/d\sqrt{d})$. This concludes the proof. ∎
D.3.2 Proof of Lemma D.2
Proof of Lemma D.2.
Fix a round $t$, an assortment $\tilde{S}_{t}$, and $\tilde{U}_{t}$. Let $U=\tilde{U}_{t}$. Define $m_{j}(\tilde{S}_{t}):=\sum_{x_{U}\in\tilde{S}_{t}}\mathbf{1}\{j\in U\}/K$. Let $\{p_{i}\}_{i\in\tilde{S}_{t}\cup\{0\}}$ and $\{q_{i}\}_{i\in\tilde{S}_{t}\cup\{0\}}$ be the probabilities of choosing item $i$ under parameterizations $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup\{j\}}$, respectively. Then, we have
$\operatorname{KL}\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)=\sum_{i\in\tilde{S}_{t}\cup\{0\}}p_{i}\log\frac{p_{i}}{q_{i}}\leq\sum_{i\in\tilde{S}_{t}\cup\{0\}}p_{i}\frac{p_{i}-q_{i}}{q_{i}}\leq\sum_{i\in\tilde{S}_{t}\cup\{0\}}\frac{(p_{i}-q_{i})^{2}}{q_{i}},$
where the first inequality holds because $\log(1+x)\leq x$ for all x>1x>-1.
Let $\hat{x}=x_{U}$. Now, we separately upper bound $(p_{i}-q_{i})^{2}/q_{i}$, by analyzing the following three cases:
Case 1. The outside option, $i=0$.
For $i=0$, $q_{i}\geq\frac{v_{0}}{v_{0}+Ke}$. Thus, we have
$|p_{i}-q_{i}|$ $=\left|\frac{v_{0}}{v_{0}+\sum_{i\in\tilde{S}_{t}}\exp\left(x_{i}^{\top}\mathbf{w}_{V}\right)}-\frac{v_{0}}{v_{0}+\sum_{i\in\tilde{S}_{t}}\exp\left(x_{i}^{\top}\mathbf{w}_{V\cup\{j\}}\right)}\right|$
$=\left|\frac{v_{0}}{v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)}-\frac{v_{0}}{v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}}\right)}\right|$
$\leq\frac{v_{0}K}{(v_{0}+K/e)^{2}}\left|\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)-\exp\left(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}}\right)\right|$
$=\frac{v_{0}K}{(v_{0}+K/e)^{2}}\left|e^{\bar{c}_{1}}(\hat{x}^{\top}\mathbf{w}_{V}-\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}})\right|$
$\leq\frac{v_{0}Ke}{(v_{0}+K/e)^{2}}\left|\hat{x}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right|\leq\frac{v_{0}Ke}{(v_{0}+K/e)^{2}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon}{\sqrt{d}},$
where the third equality holds by applying the mean value theorem for the exponential function, with $\bar{c}_{1}:=(1-u)(\hat{x}^{\top}\mathbf{w}_{V})+u(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}})$ for some $u\in(0,1)$. Then, there exist an absolute constant $C_{0}$ such that
$\frac{(p_{0}-q_{0})^{2}}{q_{0}}$ $\leq\frac{v_{0}^{2}K^{2}e^{2}}{(v_{0}+K/e)^{4}}\cdot\frac{\left(m_{j}(\tilde{S}_{t})\right)^{2}\epsilon^{2}}{d}\cdot\frac{v_{0}+Ke}{v_{0}}$
$\leq C_{0}\cdot\frac{v_{0}K^{2}}{(v_{0}+K)^{3}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$ (D.6)
where the last inequality holds since $m_{j}(\tilde{S}_{t})\leq 1$.
Case 2. $i\in\tilde{S}_{t}$ and $j\notin U$.
Then, for any $i\in\tilde{S}_{t}$ corresponding to $x_{i}=\hat{x}$ and $j\notin U$, we have
$|p_{i}-q_{i}|$ $=\left|\frac{\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)}{v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)}-\frac{\exp\left(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}}\right)}{v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}}\right)}\right|=0,$
where the last equality holds because $\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)=\exp\left(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}}\right)$, given that $j\notin U$. Thus, we get
$\sum_{i\in\tilde{S}_{t},j\notin U}\frac{(p_{i}-q_{i})^{2}}{q_{i}}=0,$ (D.7)
Case 3. $i\in\tilde{S}_{t}$ and $j\in U$.
Recall that for any $i\in\tilde{S}_{t}$, $q_{i}\geq\frac{e^{-1}}{v_{0}+Ke}$. Then, for any $i\in\tilde{S}_{t}$ corresponding to $x_{i}=\hat{x}$ and $j\in U$, we have
$|p_{i}-q_{i}|$ $=\left|\frac{\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)}{v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)}-\frac{\exp\left(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}}\right)}{v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}}\right)}\right|$
$=\left|\frac{\exp\left(\bar{c}_{2}\right)}{v_{0}+K\exp\left(\bar{c}_{2}\right)}\cdot\hat{x}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)-\frac{K\exp\left(2\bar{c}_{2}\right)}{\left(v_{0}+K\exp\left(\bar{c}_{2}\right)\right)^{2}}\cdot\hat{x}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right|$
$=\frac{\exp\left(\bar{c}_{2}\right)v_{0}}{\left(v_{0}+K\exp\left(\bar{c}_{2}\right)\right)^{2}}\left|\hat{x}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right|$
$\leq\frac{v_{0}e}{(v_{0}+K/e)^{2}}\left|\hat{x}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right|\leq\frac{v_{0}e}{(v_{0}+K/e)^{2}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon}{\sqrt{d}},$
the second equality holds by applying the mean value theorem, with $\bar{c}_{2}:=(1-u)(\hat{x}^{\top}\mathbf{w}_{V})+u(\hat{x}^{\top}\mathbf{w}_{V\cup\{j\}})$ for some $u\in(0,1)$. Then, there exist an absolute constant $C_{1}$ such that
$\sum_{i\in\tilde{S}_{t},j\in U}\frac{(p_{i}-q_{i})^{2}}{q_{i}}$ $\leq Km_{j}(\tilde{S}_{t})\cdot\frac{v_{0}^{2}e^{2}}{(v_{0}+K/e)^{4}}\cdot\frac{\left(m_{j}(\tilde{S}_{t})\right)^{2}\epsilon^{2}}{d}\cdot\frac{v_{0}+Ke}{e^{-1}}$
$\leq C_{1}\cdot\frac{v_{0}^{2}K}{(v_{0}+K)^{3}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$ (D.8)
where the last inequality holds since $m_{j}(\tilde{S}_{t})\leq 1$.
Combining Equation (D.6), (D.7), and (D.8), we derive that
$\sum_{i\in\tilde{S}_{t}\cup\{0\}}\frac{(p_{i}-q_{i})^{2}}{q_{i}}$ $\leq\left(C_{0}\cdot\frac{v_{0}K^{2}}{(v_{0}+K)^{3}}+C_{1}\cdot\frac{v_{0}^{2}K}{(v_{0}+K)^{3}}\right)\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d}$
$\leq\max\{C_{0},C_{1}\}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d}$
$=C_{\operatorname{KL}}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$
where $C_{\operatorname{KL}}=\max\{C_{0},C_{1}\}$. Since $\tilde{M}_{j}=\sum_{t=1}^{T}m_{j}(\tilde{S}_{t})$ by definition, and subsequently summing over all $t=1$ to $T$, we have
$\operatorname{KL}\left(P_{V}\|Q_{V\cup\{j\}}\right)$ $=\sum_{t=1}^{T}\mathbb{E}_{V}\left[\operatorname{KL}\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)\right]$
$\leq C_{\operatorname{KL}}\cdot\frac{v_{0}K}{(v_{0}+K)^{2}}\cdot\frac{\mathbb{E}_{V}[\tilde{M}_{j}]\epsilon^{2}}{d},$
where the equality holds by the chain rule of relative entropy (cf. Exercise 14.11 of Lattimore and Szepesvári [30]). This concludes the proof. ∎
Appendix E Proof of Theorem 2
In this section, we present the proof of Theorem 2. Note that when the rewards are uniform, the revenue increases as a function of the assortment size. Therefore, maximizing the expected revenue $R_{t}(S,\mathbf{w})$ across all possible assortments $S\in\mathcal{S}$ always contains exactly $K$ items. In other words, the size of the chosen assortment $S_{t}$ and the size of the optimal assortment $S_{t}^{\star}$ both equal to $K$.
E.1 Main Proof of Theorem 2
Before presenting the proof, we introduce useful lemmas, whose proof can be found in Appendix E.2. Lemma E.1 shows the optimistic utility for the context vectors.
Lemma E.1.
Let $\alpha_{ti}=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$. If $\mathbf{w}^{\star}\in\mathcal{C}_{t}(\delta)$, then we have
$0\leq\alpha_{ti}-x_{ti}^{\top}\mathbf{w}^{\star}\leq 2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}.$
Lemma E.3 is a $K$-free elliptical potential lemma that improves upon the one presented in Lemma 10 of Perivier and Goyal [42] in terms of $K$. Lemma 10 of Perivier and Goyal [42] states: $\sum_{s=1}^{t}\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}^{\star})p_{s}(0|S_{s},\mathbf{w}^{\star})\|x_{si}\|_{H_{s}(\mathbf{w}^{\star})^{-1}}^{2}\leq 2dK\log\left(\lambda_{t+1}+\frac{2tK}{d}\right)$ and $\sum_{s=1}^{t}\max_{i\in S_{s}}\|x_{si}\|_{H_{s}(\mathbf{w}^{\star})^{-1}}^{2}\leq 2d\left(K+\frac{1}{\kappa}\right)\log\left(\lambda_{t+1}+\frac{2tK}{d}\right)$, where $H_{t}(\mathbf{w})=\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w})+\lambda_{t}\mathbf{I}_{d}$.
Lemma E.2.
Let $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$, where $\mathcal{G}_{s}(\mathbf{w})=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w})p_{s}(j|S_{s},\mathbf{w})x_{si}x_{sj}^{\top}$. Suppose $\lambda\geq 1$. Then the following statements hold true:
  1. (1)
    $\sum_{s=1}^{t}\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(0|S_{s},\mathbf{w}_{s+1})\|x_{si}\|_{H_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{t}{d\lambda}\right)$,
  2. (2)
    $\sum_{s=1}^{t}\max_{i\in S_{s}}\|x_{si}\|_{H_{s}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{t}{d\lambda}\right)$.
Moreover, we provide a tighter bound for the second derivative of the expected revenue than that presented in Lemma 12 of Perivier and Goyal [42]. Lemma 12 of Perivier and Goyal [42] states: $\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|\leq 5$.
Lemma E.3.
Define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for any $\mathbf{u}=(u_{1},\dots,u_{K})\in\mathbb{R}^{K}$, $Q(\mathbf{u})=\sum_{i=1}^{K}\frac{\exp(u_{i})}{v_{0}+\sum_{k=1}^{K}\exp(u_{k})}$. Let $p_{i}(\mathbf{u})=\frac{\exp(u_{i})}{v_{0}+\sum_{k=1}^{K}\exp(u_{k})}$. Then, for all $i\in[K]$, we have
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|\leq\begin{cases}3p_{i}(\mathbf{u})&\text{if}\,\,\,i=j,\\ 2p_{i}(\mathbf{u})p_{j}(\mathbf{u})&\text{if}\,\,\,i\neq j.\end{cases}$
Now, we are ready to prove Theorem 2.
Proof of Theorem 2.
First, we bound the regret as follows:
$\sum_{t=1}^{T}$ $R_{t}(S_{t}^{\star},\mathbf{w}^{\star})-R_{t}(S_{t},\mathbf{w}^{\star})=\sum_{t=1}^{T}\left[\sum_{i\in S_{t}^{\star}}p_{t}(i|S_{t}^{\star},\mathbf{w}^{\star})-\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\right]$
$=\sum_{t=1}^{T}\left[\frac{\sum_{i\in S_{t}^{\star}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{v_{0}+\sum_{j\in S_{t}^{\star}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}-\frac{\sum_{i\in S_{t}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{v_{0}+\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}\right]$
$\leq\sum_{t=1}^{T}\left[\frac{\sum_{i\in S_{t}^{\star}}\exp(\alpha_{ti})}{v_{0}+\sum_{j\in S_{t}^{\star}}\exp(\alpha_{tj})}-\frac{\sum_{i\in S_{t}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{v_{0}+\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}\right]$
$\leq\sum_{t=1}^{T}\left[\frac{\sum_{i\in S_{t}}\exp(\alpha_{ti})}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha_{tj})}-\frac{\sum_{i\in S_{t}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{v_{0}+\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}\right]=\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star}),$
where the first inequality holds by Lemma E.1, and the last inequality holds by the assortment selection of Algorithm 1.
Now, we define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{K})^{\top}\in\mathbb{R}^{K}$, $Q(\mathbf{u})=\sum_{i=1}^{K}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{K}\exp(u_{j})}$. Noting that $S_{t}$ always contains $K$ elements since the expected revenue is an increasing function in the uniform reward setting, we can write $S_{t}=\{i_{1},\dots,i_{K}\}$. Moreover, for all $t\geq 1$, let $\mathbf{u}_{t}=(u_{ti_{1}},\dots u_{ti_{K}})^{\top}=(\alpha_{ti_{1}},\dots,\alpha_{ti_{K}})^{\top}$ and $\mathbf{u}^{\star}_{t}=(u_{ti_{1}}^{\star},\dots u_{ti_{K}}^{\star})^{\top}=(x_{ti_{1}}^{\top}\mathbf{w}^{\star},\dots,x_{ti_{K}}^{\top}\mathbf{w}^{\star})^{\top}$. Then, by a second order Taylor expansion, we have
$\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})$ $=\sum_{t=1}^{T}Q(\mathbf{u}_{t})-Q(\mathbf{u}^{\star}_{t})$
$=\underbrace{\sum_{t=1}^{T}\nabla Q(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(A)}}+\underbrace{\frac{1}{2}\sum_{t=1}^{T}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})^{\top}\nabla^{2}Q(\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(B)}},$ (E.1)
where $\bar{\mathbf{u}}_{t}=(\bar{u}_{ti_{1}},\dots,\bar{u}_{ti_{K}})^{\top}\in\mathbb{R}^{K}$ is the convex combination of $\mathbf{u}_{t}$ and $\mathbf{u}^{\star}_{t}$.
First, we bound the term (A).
$\sum_{t=1}^{T}\nabla Q(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}\frac{\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star})}(u_{ti}-u_{ti}^{\star})-\sum_{i\in S_{t}}\sum_{j\in S_{t}}\frac{\exp(x_{ti}^{\top}\mathbf{w}^{\star})\exp(x_{tj}^{\top}\mathbf{w}^{\star})}{(v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star}))^{2}}(u_{ti}-u_{ti}^{\star})$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})(u_{ti}-u_{ti}^{\star})-\sum_{i\in S_{t}}\sum_{j\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(j|S_{t},\mathbf{w}^{\star})(u_{ti}-u_{ti}^{\star})$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(1-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})\right)(u_{ti}-u_{ti}^{\star})$
$=\sum_{t=1}^{T}\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})(u_{ti}-u_{ti}^{\star})$
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}},$ (E.2)
where the first inequality holds by Lemma E.1, and the last inequality holds because $\beta_{t}(\delta)$ is increasing for $t\in[T]$.
Now we bound the term (B). Let $p_{i}(\bar{\mathbf{u}}_{t})=\frac{\exp(\bar{u}_{ti})}{v_{0}+\sum_{k=1}^{K}\exp(\bar{u}_{tk})}$. Then, we have
$\frac{1}{2}\sum_{t=1}^{T}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})^{\top}\nabla^{2}Q(\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})$
$=\frac{1}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t}}(u_{ti}-u_{ti}^{\star})\frac{\partial^{2}Q}{\partial i\partial j}(u_{tj}-u_{tj}^{\star})$
$=\frac{1}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t},j\neq i}(u_{ti}-u_{ti}^{\star})\frac{\partial^{2}Q}{\partial i\partial j}(u_{tj}-u_{tj}^{\star})+\frac{1}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})\frac{\partial^{2}Q}{\partial i\partial i}(u_{ti}-u_{ti}^{\star})$
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t},j\neq i}|u_{ti}-u_{ti}^{\star}|p_{i}(\bar{\mathbf{u}}_{t})p_{j}(\bar{\mathbf{u}}_{t})|u_{tj}-u_{tj}^{\star}|+\frac{3}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t}),$ (E.3)
where the inequality is by Lemma E.3. To bound the first term in Equation (E.3), by applying the AM-GM inequality, we get
$\sum_{t=1}^{T}$ $\sum_{i\in S_{t}}\sum_{j\in S_{t},j\neq i}|u_{ti}-u_{ti}^{\star}|p_{i}(\bar{\mathbf{u}}_{t})p_{j}(\bar{\mathbf{u}}_{t})|u_{tj}-u_{tj}^{\star}|$
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t}}|u_{ti}-u_{ti}^{\star}|p_{i}(\bar{\mathbf{u}}_{t})p_{j}(\bar{\mathbf{u}}_{t})|u_{tj}-u_{tj}^{\star}|$
$\leq\frac{1}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t})p_{j}(\bar{\mathbf{u}}_{t})+\frac{1}{2}\sum_{i\in S_{t}}\sum_{j\in S_{t}}(u_{tj}-u_{tj}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t})p_{j}(\bar{\mathbf{u}}_{t})$
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t}).$ (E.4)
By plugging Equation (E.4) into Equation (E.3), we have
$\frac{1}{2}\sum_{t=1}^{T}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})^{\top}\nabla^{2}Q(\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})$ $\leq\frac{5}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t})$
$\leq 10\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{i}(\bar{\mathbf{u}}_{t})\beta_{t}(\delta)^{2}\|x_{ti}\|_{H_{t}^{-1}}^{2}$
$\leq 10\sum_{t=1}^{T}\max_{i\in S_{t}}\beta_{t}(\delta)^{2}\|x_{ti}\|_{H_{t}^{-1}}^{2}$
$\leq 10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},$ (E.5)
where the second inequality holds by Lemma E.1. Combining the upper bound for the terms (A) and (B), with probability at least $1-\delta$, we have
$\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})$ $\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}$
$+10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}.$ (E.6)
Now, we bound each term of Equation (E.6) respectively. For the first term, we decompose it as follows:
$\sum_{t=1}^{T}\sum_{i\in S_{t}}$ $p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$+\sum_{t=1}^{T}\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\mathbf{w}^{\star})-p_{t}(i|S_{t},\mathbf{w}_{t+1})\right)p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$+\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(p_{t}(0|S_{t},\mathbf{w}^{\star})-p_{t}(0|S_{t},\mathbf{w}_{t+1})\right)\|x_{ti}\|_{H_{t}^{-1}}.$ (E.7)
To bound the first term on the right-hand side of Equation (E.7), we apply the Cauchy-Schwarz inequality.
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$\leq\sqrt{\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})}\sqrt{\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}^{2}}$
$\leq\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\sqrt{T}\sqrt{\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}^{2}}$
$\leq\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\sqrt{T\cdot 2d\log\left(1+\frac{T}{d\lambda}\right)},$ (E.8)
where the last inequality holds by Lemma E.2.
Now, we bound the second term on the right-hand side of Equation (E.7). Let the virtual context for the outside option be $x_{t0}=\mathbf{0}$. Then, by the mean value theorem, there exists $\bm{\xi}_{t}=(1-c)\mathbf{w}^{\star}+c\mathbf{w}_{t+1}$ for some $c\in(0,1)$ such that
$\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\mathbf{w}^{\star})-p_{t}(i|S_{t},\mathbf{w}_{t+1})\right)p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$=\sum_{i\in S_{t}}\nabla p_{t}(i|S_{t},\bm{\xi}_{t})^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$=\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\bm{\xi}_{t})x_{ti}-p_{t}(i|S_{t},\bm{\xi}_{t})\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})x_{tj}\right)^{\top}\!\!(\mathbf{w}^{\star}-\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$\leq\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\left|x_{ti}^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\right|p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$+\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left|x_{tj}^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\right|p_{t}(0|S_{t},\mathbf{w}_{t+1})$
$\leq\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}+\left(\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}\right)^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}.$
Then, since $x_{t0}=\mathbf{0}$, we can further bound the right-hand side as:
$\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}+\left(\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}\right)^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}$
$=\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}+\left(\sum_{i\in S_{t}\cup\{0\}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}\right)^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}$
$\leq\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}+\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}$
$\leq 2\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}$
$\leq 2\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}\leq 2\beta_{t}(\delta)\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},$
where the first inequality holds due to Jensen’s inequality and the second-to-last inequality holds by Lemma 1. Hence, we get
$\sum_{t=1}^{T}\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\mathbf{w}^{\star})-p_{t}(i|S_{t},\mathbf{w}_{t+1})\right)p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$ $\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}$
$\leq\frac{4d}{\kappa}\beta_{T}(\delta)\log\left(1+\frac{T}{d\lambda}\right)$ (E.9)
where the last inequality holds by Lemma E.2.
Finally, we bound the third term on the right-hand side of Equation (E.7). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
$\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(p_{t}(0|S_{t},\mathbf{w}^{\star})-p_{t}(0|S_{t},\mathbf{w}_{t+1})\right)\|x_{ti}\|_{H_{t}^{-1}}$
$=\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\nabla p_{t}(0|S_{t},\bm{\xi}^{\prime}_{t})^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$=-\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\bm{\xi}^{\prime}_{t})\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}^{\prime}_{t})x_{tj}^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$\leq\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}p_{t}(0|S_{t},\bm{\xi}^{\prime}_{t})\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}^{\prime}_{t})\|x_{tj}\|_{H_{t}^{-1}}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}$
$\leq\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}^{\prime}_{t})\|x_{tj}\|_{H_{t}^{-1}}\|\mathbf{w}^{\star}-\mathbf{w}_{t+1}\|_{H_{t}}$
$\leq\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}^{\prime}_{t})\|x_{tj}\|_{H_{t}^{-1}}$
$\leq\beta_{t}(\delta)\left(\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}\right)^{2}=\beta_{t}(\delta)\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},$
where the third inequality holds by Lemma 1, and the last inequality holds since $(\max_{i}a_{i})^{2}=\max_{i}a_{i}^{2}$ for any $a_{i}\geq 0$. Therefore, we have
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(p_{t}(0|S_{t},\mathbf{w}^{\star})-p_{t}(0|S_{t},\mathbf{w}_{t+1})\right)\|x_{ti}\|_{H_{t}^{-1}}$ $\leq\beta_{T}(\delta)\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}$
$\leq\frac{2d}{\kappa}\beta_{T}(\delta)\log\left(1+\frac{T}{d\lambda}\right),$ (E.10)
where the last inequality holds by Lemma E.2. By plugging Equation (E.8), (E.9), and (E.10) into Equation (E.7) and multiplying $2\beta_{T}(\delta)$, we get
$2\beta_{T}(\delta)$ $\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}$
$\leq 2\sqrt{2}\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\beta_{T}(\delta)\sqrt{dT}\sqrt{\log\left(1+\frac{T}{d\lambda}\right)}+\frac{12d}{\kappa}\beta_{T}(\delta)^{2}\log\left(1+\frac{T}{d\lambda}\right).$ (E.11)
Moreover, by applying Lemma E.2, we can directly bound the second term of Equation (E.6).
$10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}\leq 10\beta_{T}(\delta)^{2}\cdot\frac{2}{\kappa}d\log\left(1+\frac{T}{d\lambda}\right).$ (E.12)
Finally, plugging Equation (E.11) and (E.12) into Equation (E.6), we obtain
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})$ $\leq 2\sqrt{2}\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\beta_{T}(\delta)\sqrt{dT}\sqrt{\log\left(1+\frac{T}{d\lambda}\right)}+\frac{32d}{\kappa}\beta_{T}(\delta)^{2}\log\left(1+\frac{T}{d\lambda}\right)$
$=\tilde{\mathcal{O}}\left(\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T}+\frac{1}{\kappa}d^{2}\right),$
where $\beta_{T}(\delta)=\mathcal{O}\left(\sqrt{d}\log T\log K\right)$. This concludes the proof of Theorem 2. ∎
Remark E.1.
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, since $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right)$ (refer Corollary F.1), we have $\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(B^{3/2}e^{B}\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T}+\frac{1}{\kappa}B^{3}d^{2}\right).$ It’s important to note that one of our main goals is to explicitly demonstrate the regret depends on $K$ and $v_{0}$. In deriving such a result, the dependence on $e^{B}$ is unavoidable to our best knowledge. Note that for non-uniform rewards, the regret bound does not depend on $e^{B}$ (refer Remark H.1).
E.2 Proofs of Lemmas for Theorem 2
E.2.1 Proof of Lemma E.1
Proof of Lemma E.1.
Under the condition $\mathbf{w}^{\star}\in\mathcal{C}_{t}(\delta)$, we have
$\left|x_{ti}^{\top}\mathbf{w}_{t}-x_{ti}^{\top}\mathbf{w}^{\star}\right|\leq\|x_{ti}\|_{H_{t}^{-1}}\|\mathbf{w}_{t}-\mathbf{w}^{\star}\|_{H_{t}}\leq\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}},$
where the first inequality is by the Hölder’s inequality, and the last inequality holds by Lemma 1. Hence, it follows that
$\alpha_{ti}-x_{ti}^{\top}\mathbf{w}^{\star}=x_{ti}^{\top}\mathbf{w}_{t}-x_{ti}^{\top}\mathbf{w}^{\star}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}\leq 2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}.$
Moreover, from $x_{ti}^{\top}\mathbf{w}_{t}-x_{ti}^{\top}\mathbf{w}^{\star}\geq-\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$, we also have
$\alpha_{ti}-x_{ti}^{\top}\mathbf{w}^{\star}=x_{ti}^{\top}\mathbf{w}_{t}-x_{ti}^{\top}\mathbf{w}^{\star}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}\geq 0.$
This concludes the proof. ∎
E.2.2 Proof of Lemma E.2
Proof of Lemma E.2.
Since $xx^{\top}+yy^{\top}\succeq xy^{\top}+yx^{\top}$ for any $x,y\in\mathbb{R}^{d}$, it follows that
$\mathcal{G}_{s}(\mathbf{w}_{s+1})$
$=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(j|S_{s},\mathbf{w}_{s+1})x_{si}x_{sj}^{\top}$
$=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}-\frac{1}{2}\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(j|S_{s},\mathbf{w}_{s+1})(x_{si}x_{sj}^{\top}+x_{sj}x_{si}^{\top})$
$\succeq\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}-\frac{1}{2}\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(j|S_{s},\mathbf{w}_{s+1})(x_{si}x_{si}^{\top}+x_{sj}x_{sj}^{\top})$
$=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(j|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}.$
Hence, we have
$\mathcal{G}_{s}(\mathbf{w}_{s+1})$ $\succeq\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})\left(1-\sum_{j\in S_{s}}p_{s}(j|S_{s},\mathbf{w}_{s+1})\right)x_{si}x_{si}^{\top}$
$=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(0|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top},$ (E.13)
which implies that
$H_{t+1}\succeq H_{t}+\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})x_{ti}x_{ti}^{\top}.$
Then, we get
$\det\left(H_{t+1}\right)\geq\det\left(H_{t}\right)\left(1+\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{0}(i|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}^{2}\right).$
Since $\lambda\geq 1$, for all $t\geq 1$, we have $\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{0}(i|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}^{2}\leq 1$. Then, using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
$\sum_{s=1}^{t}\sum_{i\in S_{s}}$ $p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(0|S_{s},\mathbf{w}_{s+1})\|x_{si}\|_{H_{s}^{-1}}^{2}$
$\leq 2\sum_{s=1}^{t}\log\left(1+p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(0|S_{s},\mathbf{w}_{s+1})\|x_{si}\|_{H_{s}^{-1}}^{2}\right)$
$\leq 2\sum_{s=1}^{t}\log\left(\frac{\det(H_{s+1})}{\det(H_{s})}\right)$
$\leq 2d\log\left(\frac{\operatorname{tr}(H_{t+1})}{d\lambda}\right)\leq 2d\log\left(1+\frac{t}{d\lambda}\right).$
This proves the first inequality.
To establish the proof for the second inequality, we return to Equation (E.13):
$\mathcal{G}_{s}(\mathbf{w}_{s+1})$ $\succeq\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(0|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}\succeq\kappa\sum_{i\in S_{s}}x_{si}x_{si}^{\top},$
which implies that
$H_{t+1}=H_{t}+\mathcal{G}_{t}(\mathbf{w}_{t+1})\succeq H_{t}+\kappa\sum_{i\in S_{t}}x_{ti}x_{ti}^{\top}.$
Since $\lambda\geq 1$, for all $t\geq 1$, we have $\kappa\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}\leq\kappa$. We then conclude on the same way:
$\sum_{s=1}^{t}\max_{i\in S_{s}}\|x_{si}\|_{H_{s}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{s=1}^{t}\log\left(1+\kappa\max_{i\in S_{s}}\|x_{si}\|_{H_{s}^{-1}}^{2}\right)$
$\leq\frac{2}{\kappa}\sum_{s=1}^{t}\log\left(\frac{\det(H_{s+1})}{\det(H_{s})}\right)\leq\frac{2}{\kappa}d\log\left(1+\frac{t}{d\lambda}\right),$
which proves the second inequality. ∎
E.2.3 Proof of Lemma E.3
Proof of Lemma E.3.
Let $i,j\in[K]$. We first have
$\frac{\partial Q}{\partial i}=\frac{e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}$
Then, we get
$\frac{\partial^{2}Q}{\partial i\partial j}$
$=\frac{\mathbbm{1}_{i=j}e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}-\frac{\mathbbm{1}_{i=j}e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)+e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}$
$+\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)2e^{u_{j}}\left(v_{0}+\sum_{k=1}^{K}e^{u_{k}}\right)}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{4}}$
$=\frac{\mathbbm{1}_{i=j}e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}-\frac{\mathbbm{1}_{i=j}e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)\!+\!e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}+\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)2e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{3}}.$
Let $p_{i}(\mathbf{u})=\frac{e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ and $p_{0}(\mathbf{u})=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ . For $i=j$, we have
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ $=\bigg|p_{i}(\mathbf{u})-p_{i}(\mathbf{u})p_{j}(\mathbf{u})-p_{i}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-p_{i}(\mathbf{u})p_{j}(\mathbf{u})$
$+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg|$
$=\bigg|p_{i}(\mathbf{u})p_{0}(\mathbf{u})-2p_{i}(\mathbf{u})p_{j}(\mathbf{u})+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg|$
$=\bigg|p_{i}(\mathbf{u})p_{0}(\mathbf{u})-2p_{i}(\mathbf{u})p_{j}(\mathbf{u})p_{0}(\mathbf{u})\bigg|$
$\leq 3p_{i}(\mathbf{u})$
For $i\neq j$, we have
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ $=\bigg|-p_{i}(\mathbf{u})p_{j}(\mathbf{u})-p_{i}(\mathbf{u})p_{j}(\mathbf{u})+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg|$
$=\bigg|-2p_{i}(\mathbf{u})p_{j}(\mathbf{u})p_{0}(\mathbf{u})\bigg|$
$\leq 2p_{i}(\mathbf{u})p_{j}(\mathbf{u}).$
This concludes the proof. ∎
Appendix F Proof of Lemma 1
In this section, we provide the proof of Lemma 1. First, we present the main proof of Lemma 1, followed by the proof of the technical lemma utilized within the main proof.
F.1 Main Proof of Lemma 1
Proof of Lemma 1.
The proof is similar to the analysis presented in Zhang and Sugiyama [51]. However, their MNL choice model is constructed using a shared context $x_{t}$ and varying parameters across the choices $\mathbf{w}_{1}^{\star},\dots,\mathbf{w}_{K}^{\star}$, whereas our approach considers an MNL choice model that shares the parameter $\mathbf{w}^{\star}$ across the choices and has varying contexts for each item in the assortment $S$, $x_{t1},\dots x_{ti_{|S|}}$. Moreover, Zhang and Sugiyama [51] only consider a fixed assortment size, whereas we consider a more general setting where the assortment size can vary in each round $t$. We denote $K_{t}=|S_{t}|$ in the proof of Lemma 1. Note that $K_{t}\leq K$ for all $t\geq 1$.
Lemma F.1.
Let the update rule be
$\mathbf{w}_{t+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\tilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\|\mathbf{w}-\mathbf{w}_{t}\|_{H_{t}}^{2},$
where $\tilde{\ell}_{t}(\mathbf{w})=\ell_{t}(\mathbf{w}_{t})+\langle\mathbf{w}-\mathbf{w}_{t},\nabla\ell_{t}(\mathbf{w}_{t})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{t}\|^{2}_{\nabla^{2}\ell_{t}(\mathbf{w}_{t})}$ and $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Let $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda>0$. Then, we have
$\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t+1}}^{2}$ $\leq 2\eta\left(\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})\right)+4\lambda+12\sqrt{2}\eta\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}$
$-\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}.$ (F.1)
We first bound the first term in Equation (F.1). For simplicity, we define the softmax function at round $t$ $\bm{\sigma}_{t}(\mathbf{z}):\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$ as follows:
$[\bm{\sigma}_{t}(\mathbf{z})]_{i}=\frac{\exp([\mathbf{z}]_{i})}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})},\quad\forall i\in[K_{t}],$ (F.2)
where $\cdot]_{i}$ denotes $i$’th element of the input vector. We denote the probability of choosing the outside option as $\bm{\sigma}_{t}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})}$. Although $\bm{\sigma}_{t}(\mathbf{z})]_{0}$ is not the output of the softmax function $\bm{\sigma}_{t}(\mathbf{z})$, we represent it in a form similar to that in Equation (F.2) for simplicity. Then, the user choice model in Equation (1) can be equivalently expressed as $p_{t}(i|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{i}$ for all $i\in[K_{t}]$ and $p_{t}(0|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{0}$. Furthermore, the loss function in Equation (2) can also be written as $\ell(\mathbf{z}_{t},\mathbf{y}_{t})=\sum_{k=0}^{K_{t}}\mathbf{1}\left\{y_{ti}=1\right\}\cdot\log\left(\frac{1}{[\bm{\sigma}_{t}(\mathbf{z}_{t})]_{k}}\right)$.
Define a pseudo-inverse function of $\bm{\sigma}_{t}(\cdot)$ as $\bm{\sigma}_{t}^{+}:\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$, where $\bm{\sigma}_{t}^{+}(\mathbf{q})]_{i}=\log\left(q_{i}/(1-\|\mathbf{q}\|_{1})\right)$ for any $\mathbf{q}\in\{\mathbf{p}\in[0,1]^{K_{t}}\mid\|\mathbf{p}\|_{1}<1\}$. Then, inspired by the previous studies on binary logistic bandit [21], we decompose the regret into two terms by introducing an intermediate term.
$\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})=\underbrace{\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})}_{(a)}+\underbrace{\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})}_{(b)},$ (F.3)
where $\tilde{\mathbf{z}}_{s}:=\bm{\sigma}_{s}^{+}\left(\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[\bm{\sigma}_{s}\left((x_{sj}^{\top}\mathbf{w})_{j\in S_{s}}\right)\right]\right)$, and $P_{s}:=\mathcal{N}(\mathbf{w}_{s},cH_{s}^{-1})$ is the Gaussian distribution with mean $\mathbf{w}_{s}$ and covariance matrix $cH_{s}^{-1}$, where $c>0$ is a positive constant to be specified later. We first show that the term $(a)$ is bounded by $\mathcal{O}\left(\log K(\log t)^{2}\right)$ with high probability.
Lemma F.2.
Let $\delta\in(0,1]$. Under Assumptions 1, for all $t\in[T]$, with probability at least $1-\delta$, we have
$\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})$ $\leq 11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2.$
Furthermore, we can bound the term $(b)$ by the following lemma.
Lemma F.3.
For any $c>0$, let $\lambda\geq\max\{2,72cd\}$. Then, under Assumption 1, for all $t\geq 1$, we have
$\sum_{s=1}^{t}\left(\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})-\ell_{s}(\mathbf{w}_{s+1})\right)\leq\frac{1}{2c}\sum_{s=1}^{t}\|\mathbf{w}_{s}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right).$
Now, we are ready to prove the Lemma 1. By combining Lemma F.1, Lemma F.2, and Lemma F.3, we derive that
$\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t+1}}^{2}$
$\leq 2\eta\Bigg[\!11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg]$
$+\!4\lambda\!+12\sqrt{2}\eta\sum_{s=1}^{t}\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}+\left(\frac{\eta}{c}\!-\!1\right)\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$\leq 2\eta\Bigg[\!11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg]$
$+\!4\lambda\!=:\beta^{\prime}_{t+1}(\delta)^{2}=\mathcal{O}\left(d(\log t\log K)^{2}\right),$ (F.4)
where the second inequality holds because by setting $c=7\eta/6$ and $\lambda\geq\max\{84\sqrt{2}\eta,84d\eta\}$, we obtain:
$12\sqrt{2}\eta\sum_{s=1}^{t}$ $\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}+\left(\frac{\eta}{c}\!-\!1\right)\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$=12\sqrt{2}\eta\sum_{s=1}^{t}\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}-\frac{1}{7}\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$\leq\left(12\sqrt{2}\eta-\frac{\lambda}{7}\right)\sum_{s=1}^{t}\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}\leq 0,$
where the first inequality holds since $H_{s}\succeq\lambda\mathbf{I}_{d}$.
By setting $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda=84\sqrt{2}d\eta$, we derive that
$\|\mathbf{w}_{t}-\mathbf{w}^{\star}\|_{H_{t}}$ $\leq\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\|_{H_{t}}+\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t}}$
$\leq\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\|_{H_{t}}+\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t+1}}$
$\leq\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\|_{H_{t}}+\beta^{\prime}_{t+1}(\delta)$
$\leq\frac{2\eta}{\lambda}+\beta^{\prime}_{t+1}(\delta):=\beta_{t}(\delta)=\mathcal{O}\left(\sqrt{d}\log t\log K\right).$
where the third inequality follows from the definition of $\beta^{\prime}_{t+1}(\delta)$ in Equation (F.4), and the last inequality holds by Lemma F.9.
This concludes the proof of Lemma 1. ∎
Corollary F.1.
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, then $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right)$.
Proof of Corollary F.1.
When $\|\mathbf{w}\|_{2}\leq B$, following the same analysis as in Lemma 1, we can set $\eta=\frac{1}{2}\log(K+1)+B+1$, $c=7\eta/6$, and $\lambda=\mathcal{O}(dB\log K)$. Under these settings, we have
$\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t+1}}$
$\leq\sqrt{2\eta\Bigg[\!11\cdot\left(3\log(1+(K+1)t)+B+2\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg]+\!4B^{2}\lambda\!}$
$=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right).$
which concludes the proof. ∎
F.2 Proofs of Lemmas for Lemma 1
F.2.1 Proof of Lemma F.1
Proof of Lemma F.1.
Let $\tilde{\ell}_{s}(\mathbf{w})=\ell_{s}(\mathbf{w}_{s})+\langle\nabla\ell_{s}(\mathbf{w}_{s}),\mathbf{w}-\mathbf{w}_{s}\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{s}\|^{2}_{\nabla^{2}\ell_{s}(\mathbf{w}_{s})}$ be a second-order approximation of the original function $\ell_{s}(\mathbf{w})$ at the point $\mathbf{w}_{s}$. The update rule in Equation (3) can also be expressed as
$\mathbf{w}_{s+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\tilde{\ell}_{s}(\mathbf{w})+\frac{1}{2\eta}\|\mathbf{w}-\mathbf{w}_{s}\|_{H_{s}}^{2}.$
Then, by Lemma F.4, we have
$\langle\nabla\tilde{\ell}_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle\leq\frac{1}{2\eta}\left(\|\mathbf{w}_{s}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{H_{s}}^{2}\right).$ (F.5)
To utilize Lemma F.6, we can rewrite the loss function as $\ell\left((x_{si}^{\top}\mathbf{w})_{i\in S_{s}},\mathbf{y}_{s}\right)=\ell_{s}(\mathbf{w})$. Consequently, according to Lemma F.6, it follows that
$\ell_{s}(\mathbf{w}_{s+1})-\ell_{s}(\mathbf{w}^{\star})\leq\langle\nabla\ell_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle-\frac{1}{\zeta}\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|^{2}_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})},$ (F.6)
where $\zeta=\log(K+1)+4$. Then, by combining Equation (F.5) and (F.6), we have
$\ell_{s}(\mathbf{w}_{s+1})-\ell_{s}(\mathbf{w}^{\star})$ $\leq\langle\nabla\ell_{s}(\mathbf{w}_{s+1})-\nabla\tilde{\ell}_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle$
$+\frac{1}{\zeta}\left(\|\mathbf{w}_{s}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{H_{s+1}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{H_{s}}^{2}\right).$
In above, we can further bound the first term of the right-hand side as:
$\langle\nabla\ell_{s}(\mathbf{w}_{s+1})$- $\nabla\tilde{\ell}_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle$
$=\langle\nabla\ell_{s}(\mathbf{w}_{s+1})-\nabla\ell_{s}(\mathbf{w}_{s})-\nabla^{2}\ell_{s}(\mathbf{w}_{s})(\mathbf{w}_{s+1}-\mathbf{w}_{s}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle$
$=\langle D^{3}\ell_{s}(\bm{\xi}_{s+1})[\mathbf{w}_{s+1}-\mathbf{w}_{s}](\mathbf{w}_{s+1}-\mathbf{w}_{s}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle$
$\leq 3\sqrt{2}\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{2}\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{\nabla^{2}\ell_{s}(\bm{\xi}_{s+1})}^{2}$
$\leq 6\sqrt{2}\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{\nabla^{2}\ell_{s}(\bm{\xi}_{s+1})}^{2}$
$\leq 6\sqrt{2}\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{2}^{2}$
where in the second equality, we apply the Taylor expansion by introducing $\bm{\xi}_{s+1}$, a convex combination of $\mathbf{w}_{s+1}$ and $\mathbf{w}_{s}$. The first inequality follows from Lemma C.1 and Proposition C.1, the second inequality holds by Assumption 1, and the last inequality holds because
$\nabla^{2}\ell_{s}(\bm{\xi}_{s+1})$ $=\mathcal{G}_{s}(\bm{\xi}_{s+1})$
$=\sum_{i\in S_{s}}p_{s}(i|S_{s},\bm{\xi}_{s+1})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\bm{\xi}_{s+1})p_{s}(j|S_{s},\bm{\xi}_{s+1})x_{si}x_{sj}^{\top}$
$=\!\!\sum_{i\in S_{s}\!\cup\{0\}}p_{s}(i|S_{s},\bm{\xi}_{s+1})x_{si}x_{si}^{\top}-\!\!\sum_{i\in S_{s}\!\cup\{0\}}\sum_{j\in S_{s}\!\cup\{0\}}p_{s}(i|S_{s},\bm{\xi}_{s+1})p_{s}(j|S_{s},\bm{\xi}_{s+1})x_{si}x_{sj}^{\top}$
$=\mathbb{E}_{i\sim p_{s}(\cdot|S_{,}\bm{\xi}_{s+1})}\left[x_{si}x_{si}^{\top}\right]-\mathbb{E}_{i\sim p_{s}(\cdot|S_{s},\bm{\xi}_{s+1})}\left[x_{si}\right]\left(\mathbb{E}_{i\sim p_{s}(\cdot|S_{s},\bm{\xi}_{s+1})}\left[x_{si}\right]\right)^{\top}$
$\preceq\mathbb{E}_{i\sim p_{s}(\cdot|S_{s},\bm{\xi}_{s+1})}\left[x_{si}x_{si}^{\top}\right]\preceq\mathbf{I}_{d},$
where the third equality holds by setting $x_{s0}=\mathbf{0}$ for all $s\geq 1$.
Now, by taking the summation over $s$ and rearranging the terms, we obtain
$\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t+1}}^{2}$
$\leq\zeta\left(\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})\right)+\|\mathbf{w}_{1}-\mathbf{w}^{\star}\|_{H_{1}}^{2}+6\sqrt{2}\zeta\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}$
$-\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$\leq\zeta\left(\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})\right)+4\lambda+6\sqrt{2}\zeta\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}-\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2},$
where the last inequality holds since $\|\mathbf{w}_{1}-\mathbf{w}^{\star}\|_{H_{1}}^{2}\leq\lambda\|\mathbf{w}_{1}-\mathbf{w}^{\star}\|_{2}^{2}\leq 4\lambda$. Plugging in $\zeta=2\eta$, we conclude the proof. ∎
F.2.2 Proof of Lemma F.2
Proof of Lemma F.2.
Since the norm of $\tilde{\mathbf{z}}_{s}=\bm{\sigma}_{s}^{+}\left(\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[\bm{\sigma}_{s}\left((x_{sj}^{\top}\mathbf{w})_{j\in S_{s}}\right)\right]\right)$ is unbounded in general, as suggested by Foster et al. [23], we use the smoothed version $\tilde{\mathbf{z}}^{\mu}_{s}=\bm{\sigma}_{s}^{+}\left(\operatorname{smooth}^{\mu}_{s}\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[\bm{\sigma}_{s}\left((x_{sj}^{\top}\mathbf{w})_{j\in S_{s}}\right)\right]\right)$ as an intermediate-term, where the smooth function is defined by $\operatorname{smooth}^{\mu}_{s}(\mathbf{q})=(1-\mu)\mathbf{q}+\mu\mathbf{1}/(K_{s}+1)$, where $\mathbf{1}\in\mathbb{R}^{K_{s}}$ is an all one vector.
Note that $\tilde{\mathbf{z}}_{s}^{\mu}=\bm{\sigma}_{s}^{+}(\operatorname{smooth}^{\mu}_{s}(\bm{\sigma}_{s}(\tilde{\mathbf{z}}_{s})))$ by the definition of the pseudo inverse function $\bm{\sigma}_{s}^{+}$ such that $\bm{\sigma}_{s}^{+}(\bm{\sigma}_{s}(\mathbf{q}))=\mathbf{q}$ for any $\mathbf{q}\in\{\mathbf{p}\in[0,1]^{K_{s}}\mid\|\mathbf{p}\|_{1}<1\}$. Then, by Lemma F.7, we have
$\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s}^{\mu},\mathbf{y}_{s})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq 2\mu t,\quad\text{and}\quad\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu).$ (F.7)
Hence, to prove the lemma, we need only to bound the gap between the loss of $\mathbf{w}^{\star}$ and $\tilde{\mathbf{z}}_{s}^{\mu}$. To enhance clarity in our presentation, let $\ell(\mathbf{z}_{s}^{\star},\mathbf{y}_{s})=\ell_{s}(\mathbf{w}^{\star})$, where $\mathbf{z}_{s}^{\star}=\left(x_{sj}^{\top}\mathbf{w}^{\star}\right)_{j\in S_{s}}\in\mathbb{R}^{K_{s}}$. Then, we have
$\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s}^{\mu},\mathbf{y}_{s})$ $=\sum_{s=1}^{t}\ell(\mathbf{z}_{s}^{\star},\mathbf{y}_{s})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s}^{\mu},\mathbf{y}_{s})$
$\leq\sum_{s=1}^{t}\langle\nabla_{z}\ell(\mathbf{z}_{s}^{\star},\mathbf{y}_{s}),\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\rangle-\sum_{s=1}^{t}\frac{1}{c_{\mu}}\|\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\|_{\nabla_{z}^{2}\ell(\mathbf{z}_{s}^{\star},\mathbf{y}_{s})}^{2}$
$=\sum_{s=1}^{t}\langle\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s},\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\rangle-\sum_{s=1}^{t}\frac{1}{c_{\mu}}\|\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2},$ (F.8)
where $c_{\mu}=\log(K+1)+2\log(1+(K+1)/\mu)+2$, the inequality holds by Lemma F.6, and the last equality holds by a direct calculation of the first order and Hessian of the logistic loss as follows:
$\nabla_{z}\ell(\mathbf{z}_{s},\mathbf{y}_{s})=\bm{\sigma}_{s}(\mathbf{z}_{s})-\mathbf{y}_{s},\quad\nabla_{z}^{2}\ell(\mathbf{z}_{s},\mathbf{y}_{s})=\operatorname{diag}(\bm{\sigma}_{s}(\mathbf{z}_{s}))-\bm{\sigma}_{s}(\mathbf{z}_{s})\bm{\sigma}_{s}(\mathbf{z}_{s})^{\top}.$
We first bound the first term of the right-hand side. Define $\mathbf{d}_{s}=(\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu})/(c_{\mu}+1)$. Let $\mathbf{d}^{\prime}_{s}$ be $\mathbf{d}_{s}$ extended with zero padding. Specifically, we define $\mathbf{d}^{\prime}_{s}=[\mathbf{d}_{s}^{\top},0,\dots,0]^{\top}\in\mathbb{R}^{K}$, where the zeros are appended to increase the dimension of $\mathbf{d}_{s}$ to $K$. Similarly, we also extend $\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s}$ with zero padding and define $\bm{\varepsilon}_{s}=[(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})^{\top},0,\dots,0]^{\top}\in\mathbb{R}^{K}$.
Then, one can easily verify that $\|\mathbf{d}^{\prime}_{s}\|_{\infty}\leq 1$ since $\|\mathbf{z}_{s}^{\star}\|_{\infty}\leq\max_{i\in S_{s}}\|x_{si}\|_{2}\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu)$. On the other hand, $\mathbf{d}_{s}^{\prime}$ is $\mathcal{F}_{s}$-measurable since $\mathbf{z}_{s}^{\star}$ and $\tilde{\mathbf{z}}_{s}^{\mu}$ are independent of $\mathbf{y}_{s}$. Moreover, we have $\|\mathbf{d}^{\prime}_{s}\|^{2}_{\mathbb{E}[\bm{\varepsilon}_{s}\bm{\varepsilon}_{s}^{\top}\mid\mathcal{F}_{s}]}=\|\mathbf{d}_{s}\|^{2}_{\mathbb{E}\left[(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})^{\top}\mid\mathcal{F}_{s}\right]}=\|\mathbf{d}_{s}\|^{2}_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}$ and $\|\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s}\|_{1}\leq 2$. Thus, by Lemma F.5, with probability at least $1-\delta$, for any $t\geq 1$, we have
$\sum_{s=1}^{t}\langle\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s},\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\rangle=(c_{\mu}+1)\sum_{s=1}^{t}\langle\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s},\mathbf{d}_{s}\rangle$
$=(c_{\mu}+1)\sum_{s=1}^{t}\langle\bm{\varepsilon}_{s},\mathbf{d}^{\prime}_{s}\rangle$
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\cdot\left(\frac{2\sqrt{1+\frac{1}{\tilde{\lambda}}\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}}{\delta}\right)\right)$
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right),$ (F.9)
where in the second inequality, we set $\tilde{\lambda}>1$, and the last inequality holds because $\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}=\mathbf{d}_{s}^{\top}\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})\mathbf{d}_{s}\leq 2$. Then, combining Equation (F.8) and (F.9), we obtain
$\sum_{s=1}^{t}$ $\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s}^{\mu},\mathbf{y}_{s})$
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)-\sum_{s=1}^{t}\frac{1}{c_{\mu}}\|\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}$
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)-(c_{\mu}+1)\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}$
$\leq(c_{\mu}+1)\left(\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}\right)+(c_{\mu}+1)\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}$
$-(c_{\mu}+1)\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}$
$=(c_{\mu}+1)\left(\frac{17}{16}\tilde{\lambda}+2\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+\frac{16}{\tilde{\lambda}}\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right),$ (F.10)
where the third inequality holds due to the AM-GM inequality. Finally, combining Equation (F.7) and (F.10), by setting $\mu=1/t$ and $\tilde{\lambda}=\frac{16}{\sqrt{17}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)$, we have
$\sum_{s=1}^{t}\left(\ell_{s}(\mathbf{w}^{\star})-\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\right)$ $\leq(c_{\mu}+1)(2\sqrt{17}+2)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2\mu t$
$\leq 11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2$
where the last inequality holds by the definition of $c_{\mu}=\log(K+1)+2\log(1+(K+1)/\mu)+2$. This concludes the proof. ∎
F.2.3 Proof of Lemma F.3
Proof of Lemma F.3.
The proof with an observation from Proposition 2 in Foster et al. [23], which notes that $\tilde{\mathbf{z}}_{s}$ is an aggregation forecaster for the logistic function. Hence, it satisfies
$\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq-\log\left(\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[e^{-\ell_{s}(\mathbf{w})}\right]\right)=-\log\left(\frac{1}{Z_{s}}\int_{\mathbb{R}^{d}}e^{-L_{s}(\mathbf{w})}\mathrm{d}\mathbf{w}\right),$ (F.11)
where $L_{s}(\mathbf{w}):=\ell_{s}(\mathbf{w})+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s}\|_{H_{s}}^{2}$ and $Z_{s}:=\sqrt{(2\pi)^{d}c|H_{s}^{-1}|}$.
Then, by the quadratic approximation, we get
$\tilde{L}_{s}(\mathbf{w})=L_{s}(\mathbf{w}_{s+1})+\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}.$ (F.12)
Applying Lemma F.8 and considering the fact that $\ell_{s}$ is $3\sqrt{2}$-self-concordant-like function by Proposition C.1, we have
$L_{s}(\mathbf{w})\leq\tilde{L}_{s}(\mathbf{w})+e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}.$ (F.13)
We define the function $\tilde{f}_{s+1}:\mathcal{W}\rightarrow\mathbb{R}$ as
$\tilde{f}_{s+1}(\mathbf{w})=\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}-e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right).$
Then, we can establish a lower bound for the expectation in Equation (F.11) as follows:
$\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[e^{-\ell_{s}(\mathbf{w})}\right]$ $=\frac{1}{Z_{s}}\int_{\mathbb{R}^{d}}\exp(-L_{s}(\mathbf{w}))\mathrm{d}\mathbf{w}$
$\geq\frac{1}{Z_{s}}\int_{\mathbb{R}^{d}}\exp(-\tilde{L}_{s}(\mathbf{w})-e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla\ell_{s}(\mathbf{w}_{s+1})}^{2})\mathrm{d}\mathbf{w}$
$=\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))}{Z_{s}}\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\cdot\exp(-\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle)\mathrm{d}\mathbf{w},$ (F.14)
where the first inequality holds by Equation (F.13) and the last equality holds by Equation (F.12). We define $\tilde{Z}_{s+1}=\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\mathrm{d}\mathbf{w}\leq+\infty$. Moreover, we denote the distribution whose density function is $\tilde{f}_{s+1}(\mathbf{w})/\tilde{Z}_{s+1}$ as $\tilde{P}_{s+1}$. Then, we can rewrite Equation (F.14) as follows:
$\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[e^{-\ell_{s}(\mathbf{w})}\right]$ $\geq\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))\tilde{Z}_{s+1}}{Z_{s}}\mathbb{E}_{\mathbf{w}\sim\tilde{P}_{s+1}}\left[\exp(-\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle)\right]$
$\geq\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))\tilde{Z}_{s+1}}{Z_{s}}\exp\left(-\mathbb{E}_{\mathbf{w}\sim\tilde{P}_{s+1}}\left[\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle\right]\right)$
$=\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))\tilde{Z}_{s+1}}{Z_{s}},$ (F.15)
where the second inequality follows from Jensen’s inequality, and the equality holds because $\tilde{P}_{s+1}$ is symmetric around $\mathbf{w}_{s+1}$, thus $\mathbb{E}_{\mathbf{w}\sim\tilde{P}_{s+1}}\left[\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle\right]=0$.
By plugging Equation (F.15) into Equation (F.11), we have
$\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq L_{s}(\mathbf{w}_{s+1})+\log Z_{s}-\log\tilde{Z}_{s+1}.$ (F.16)
In the above, we can bound the last term, $\log\tilde{Z}_{s+1}$, by
$-\log\tilde{Z}_{s+1}$ $=-\log\left(\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}-e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right)\mathrm{d}\mathbf{w}\right)$
$=-\log\left(\widehat{Z}_{s+1}\cdot\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\exp\left(-e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right)\right]\right)$
$\leq-\log\widehat{Z}_{s+1}+\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right]$
$=-\log Z_{s}+\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right],$ (F.17)
where $\widehat{P}_{s+1}=\mathcal{N}(\mathbf{w}_{s+1},cH_{s}^{-1})$ and $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}$. In Equation (F.17), the inequality holds due to Jensen’s inequality, and the last inequality is by the fact that $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}=\sqrt{(2\pi)^{d}c|H_{s}^{-1}|}=Z_{s}$.
By applying the Cauchy-Schwarz inequality, we can further bound the second term on the right-hand side of Equation (F.17) by
$\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}$ $\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right]$
$\leq\underbrace{\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{36\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\right]}}_{\texttt{(a)-1}}\underbrace{\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}}_{\texttt{(a)-2}}.$ (F.18)
Note that, since $\widehat{P}_{s+1}=\mathcal{N}(\mathbf{w}_{s+1},cH_{s}^{-1})$, there exist orthogonal bases $\mathbf{e}_{1},\dots,\mathbf{e}_{d}\in\mathbb{R}^{d}$ such that $\mathbf{w}-\mathbf{w}_{s+1}$ follows the same distribution as
$\sum_{j=1}^{d}\sqrt{c\lambda_{j}\left(H_{s}^{-1}\right)}X_{j}\mathbf{e}_{j},\quad\text{where }X_{j}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,1),\forall j\in[d],$ (F.19)
and $\lambda_{j}\left(H_{s}^{-1}\right)$ denotes the $j$-th largest eigenvalue of $H_{s}^{-1}$. Then, we can bound the term (a)-1 in Equation (F.18) as follows:
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{36\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\right]}$ $=\sqrt{\mathbb{E}_{X_{j}}\left[\prod_{j=1}^{d}e^{36c\lambda_{j}\left(H_{s}^{-1}\right)X_{j}^{2}}\right]}\leq\sqrt{\prod_{j=1}^{d}\mathbb{E}_{X_{j}}\left[e^{\frac{36c}{\lambda}X_{j}^{2}}\right]}$
$=\left(\mathbb{E}_{X\sim\chi^{2}}\left[e^{\frac{36c}{\lambda}X}\right]\right)^{\frac{d}{2}}\leq\mathbb{E}_{X\sim\chi^{2}}\left[e^{\frac{18cd}{\lambda}X}\right],$
where the first inequality holds since $\lambda_{j}\left(H_{s}^{-1}\right)\leq\frac{1}{\lambda}$. In the second equality, $\chi^{2}$ denotes the chi-square distribution, and the last inequality is due to Jensen’s inequality. By setting $\lambda\geq 72cd$, we get
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{36\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\right]}\leq\mathbb{E}_{X\sim\chi^{2}}\left[e^{\frac{X}{4}}\right]\leq\sqrt{2},$ (F.20)
where the last inequality holds due to the fact that the moment-generating function for $\chi^{2}$-distribution is bounded by $\mathbb{E}_{X\sim\chi^{2}}[e^{tX}]\leq 1/\sqrt{1-2t}$ for all $t\leq 1/2$.
Now, we bound the term (a)-2 in Equation (F.18).
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}$ $=\sqrt{\mathbb{E}_{\mathbf{w}\sim\mathcal{N}(0,cH_{s}^{-1})}\left[\|\mathbf{w}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}$
$=\sqrt{\mathbb{E}_{\mathbf{w}\sim\mathcal{N}(0,c\bar{H}_{s}^{-1})}\left[\|\mathbf{w}\|_{2}^{4}\right]},$
where $\bar{H}_{s}=(\nabla^{2}\ell_{s}(\mathbf{w}_{s+1}))^{-1/2}H_{s}(\nabla^{2}\ell_{s}(\mathbf{w}_{s+1}))^{-1/2}$. Let $\bar{\lambda}_{j}=\lambda_{j}\left(c\bar{H}_{s}^{-1}\right)$ be the $j$-th largest eigenvalue of the matrix $c\bar{H}_{s}^{-1}$. Then, conducting an analysis similar to that in Equation (F.19) yields that
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\mathcal{N}(0,c\bar{H}_{s}^{-1})}\left[\|\mathbf{w}\|_{2}^{4}\right]}$ $=\sqrt{\mathbb{E}_{X_{j}\sim\mathcal{N}(0,1)}\left[\left\|\sum_{j=1}^{d}\sqrt{\bar{\lambda}_{j}}X_{j}\mathbf{e}_{j}\right\|_{2}^{4}\right]}$
$=\sqrt{\mathbb{E}_{X_{j}\sim\mathcal{N}(0,1)}\left[\left(\sum_{j=1}^{d}\bar{\lambda}_{j}X_{j}^{2}\right)^{2}\right]}$
$=\sqrt{\sum_{j=1}^{d}\sum_{j^{\prime}=1}^{d}\bar{\lambda}_{j}\bar{\lambda}_{j^{\prime}}\mathbb{E}_{X_{j},X_{j^{\prime}}\sim\mathcal{N}(0,1)}\left[X_{j}^{2}X_{j^{\prime}}^{2}\right]}$
$\leq\sqrt{3\sum_{j=1}^{d}\sum_{j^{\prime}=1}^{d}\bar{\lambda}_{j}\bar{\lambda}_{j^{\prime}}}=\sqrt{3}c\operatorname{Tr}\left(\bar{H}_{s}^{-1}\right),$
where the inequality holds due to $\mathbb{E}_{X_{j},X_{j^{\prime}}\sim\mathcal{N}(0,1)}[X_{j}^{2}X_{j^{\prime}}^{2}]\leq 3$ for all $j,j^{\prime}\in[d]$, and the last equality holds because $\sum_{j=1}^{d}\bar{\lambda}_{j}=\operatorname{Tr}\left(c\bar{H}_{s}^{-1}\right)$. Here, $\operatorname{Tr}(A)$ denotes the trace of the matrix $A$.
We define the matrix $M_{s+1}:=\lambda\mathbf{I}_{d}/2+\sum_{\tau=1}^{s}\nabla^{2}\ell_{\tau}(\mathbf{w}_{\tau+1})$. Under the condition $\lambda\geq 2$, for any $s\in[T]$ and $\mathbf{w}\in\mathcal{W}$, we have $\nabla^{2}\ell_{s}(\mathbf{w})\preceq\mathbf{I}_{d}\leq\frac{\lambda}{2}\mathbf{I}_{d}$. Thus, we have $H_{s}\succeq M_{s+1}$. Then, we can bound the trace as follows:
$\operatorname{Tr}\left(\bar{H}_{s}^{-1}\right)$ $=\operatorname{Tr}\left(H_{s}^{-1}\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})\right)\leq\operatorname{Tr}\left(M_{s+1}^{-1}\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})\right)$
$=\operatorname{Tr}\left(M_{s+1}^{-1}(M_{s+1}-M_{s})\right)\leq\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})},$
where the last inequality holds by Lemma 4.5 of Hazan et al. [24]. Therefore we can bound the term (a)-2 as
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}\leq\sqrt{3}c\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}.$ (F.21)
By plugging Equation (F.20) and (F.21) into Equation (F.18), we have
$\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right]\leq\sqrt{6}c\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}.$ (F.22)
Combining Equation (F.16), (F.17), and (F.22), and taking summation over $s$, we derive that
$\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})$ $\leq\sum_{s=1}^{t}L_{s}(\mathbf{w}_{s+1})+\sqrt{6}c\sum_{s=1}^{t}\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}$
$=\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})+\frac{1}{2c}\sum_{s=1}^{t}\|\mathbf{w}_{s}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}+\sqrt{6}c\sum_{s=1}^{t}\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}$
$=\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})+\frac{1}{2c}\sum_{s=1}^{t}\|\mathbf{w}_{s}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}+\sqrt{6}c\log\left(\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}\left(\frac{\lambda}{2}\mathbf{I}_{d}\right)}\right)$
$\leq\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})+\frac{1}{2c}\sum_{s=1}^{t}\|\mathbf{w}_{s}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}+\sqrt{6}c\cdot d\log\left(1+\frac{t+1}{2\lambda}\right),$
By rearranging the terms, we conclude the proof. ∎
F.3 Technical Lemmas for Lemma 1
Lemma F.4 (Proposition 4.1 of Campolongo and Orabona 10).
Let the $\mathbf{w}_{t+1}$ be the solution of the update rule
$\mathbf{w}_{t+1}=\arg\min_{\mathbf{w}\in\mathcal{V}}\eta_{t}\ell_{t}(\mathbf{w})+D_{\psi}(\mathbf{w},\mathbf{w}_{t}),$
where $\mathcal{V}\subseteq\mathcal{W}\subseteq\mathbb{R}^{d}$ is a non-empty convex set and $D_{\psi}(\mathbf{w}_{1},\mathbf{w}_{2})=\psi(\mathbf{w}_{1})-\psi(\mathbf{w}_{2})-\langle\nabla\psi(\mathbf{w}_{2}),\mathbf{w}_{1}-\mathbf{w}_{2}\rangle$ is the Bregman Divergence w.r.t. a strictly convex and continuously differentiable function $\psi:\mathcal{W}\rightarrow\mathbb{R}$. Further supposing $\psi(\mathbf{w})$ is $1$-strongly convex w.r.t. a certain norm $\|\cdot\|$ in $\mathcal{W}$, then there exists a $\mathbf{g}^{\prime}_{t}\in\partial\ell_{t}(\mathbf{w}_{t+1})$ such that
$\langle\eta_{t}\mathbf{g}^{\prime}_{t},\mathbf{w}_{t+1}-\mathbf{u}\rangle\leq\langle\nabla\psi(\mathbf{w}_{t})-\nabla\psi(\mathbf{w}_{t+1}),\mathbf{w}_{t+1}-\mathbf{u}\rangle$
for any $\mathbf{u}\in\mathcal{W}$.
Lemma F.5 (Lemma 15 of Zhang and Sugiyama 51).
Let $\{\mathcal{F}_{t}\}_{t=1}^{\infty}$ be a filtration. Let $\{\mathbf{z}_{t}\}_{t=1}^{\infty}$ be a stochastic process in $\mathcal{B}_{2}(K)=\{\mathbf{z}\in\mathbb{R}^{K}\mid\|\mathbf{z}\|_{\infty}\leq 1\}$ such that $\mathbf{z}_{t}$ is $\mathcal{F}_{t}$ measurable. Let $\{\bm{\bm{\varepsilon}}_{t}\}_{t=1}^{\infty}$ be a martingale difference sequence such that $\bm{\varepsilon}_{t}\in\mathbb{R}^{K}$ is $\mathcal{F}_{t+1}$ measurable. Furthermore, assume that, conditional on $\mathcal{F}_{t}$, we have $\|\bm{\varepsilon}_{t}\|_{1}\leq 2$ almost surely. Let $\Sigma_{t}=\mathbb{E}[\bm{\varepsilon}_{t}\bm{\varepsilon}_{t}^{\top}|\mathcal{F}_{t}]$. and $\lambda>0$. Then, for any $t\geq 1$ define
$U_{t}=\sum_{s=1}^{t-1}\langle\bm{\varepsilon}_{s},\mathbf{z}_{s}\rangle\quad\text{and}\quad H_{t}=\lambda+\sum_{s=1}^{t-1}\|\mathbf{z}_{s}\|_{\Sigma_{s}}^{2},$
Then, for any $\delta\in(0,1]$, we have
$\Pr\left[\exists t\geq 1,U_{t}\geq\sqrt{H_{t}}\left(\frac{\sqrt{\lambda}}{4}+\frac{4}{\sqrt{\lambda}}\log\left(\sqrt{\frac{H_{t}}{\lambda}}\right)+\frac{4}{\sqrt{\lambda}}\log\left(\frac{2}{\delta}\right)\right)\right]\leq\delta.$
Lemma F.6 (Lemma 1 of Zhang and Sugiyama 51).
Let $C>0$, $\mathbf{a}\in[-C,C]^{K}$, $\mathbf{y}\in\mathbb{R}^{K+1}$ be a one-hot vector and $\mathbf{b}\in\mathbb{R}^{K}$. Then, we have
$\ell(\mathbf{a},\mathbf{y})\geq\ell(\mathbf{b},\mathbf{y})+\nabla\ell(\mathbf{b},\mathbf{y})^{\top}(\mathbf{a}-\mathbf{b})+\frac{1}{\log(K+1)+2(C+1)}(\mathbf{a}-\mathbf{b})^{\top}\nabla^{2}\ell(\mathbf{b},\mathbf{y})(\mathbf{a}-\mathbf{b}).$
Lemma F.7 (Lemma 17 of Zhang and Sugiyama 51).
Let $\mathbf{z}\in\mathbb{R}^{K}$ be a $K$-dimensional vector. Let $\ell(\mathbf{z},\mathbf{y})=\sum_{k=0}^{K}\mathbf{1}\{y_{i}=1\}\cdot\log\left(\frac{1}{[\bm{\sigma}(\mathbf{z})]_{k}}\right)$, where $\mathbf{y}=[y_{0},\dots,y_{K}]^{\top}\in\mathbb{R}^{K+1}$, and the softmax function $\bm{\sigma}(\mathbf{z}):\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ is defined as $\bm{\sigma}(\mathbf{z})]_{i}=\frac{\exp([\mathbf{z}]_{i})}{v_{0}+\sum_{k=1}^{K}\exp([\mathbf{z}]_{k})}$ for all $i\in[K]$, and $\bm{\sigma}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K}\exp([\mathbf{z}]_{k})}$. Define $\mathbf{z}^{\mu}:=\bm{\sigma}^{+}\left(\operatorname{smooth}_{\mu}(\bm{\sigma}(\mathbf{z}))\right)$, where $\operatorname{smooth}_{\mu}(\mathbf{q})=(1-\mu)\mathbf{q}+\mu\mathbf{1}/(K+1)$. Then, for $\mu\in[0,1/2]$, we have
$\ell(\mathbf{z}^{\mu},\mathbf{y})-\ell(\mathbf{z},\mathbf{y})\leq 2\mu$
We also have $\|\mathbf{z}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu)$.
Lemma F.8 (Lemma 18 of Zhang and Sugiyama 51).
Let $L_{t}(\mathbf{w})=\ell_{t}(\mathbf{w})+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{t}\|_{H_{t}}^{2}$. Assume that $\ell_{t}$ is a $M$-self-concordant-like function. Then, for any $\mathbf{w},\mathbf{w}_{t}\in\mathcal{W}$, the quadratic approximation $\tilde{L}_{t}(\mathbf{w})=L_{t}(\mathbf{w}_{t+1})+\langle\nabla^{2}L_{t}(\mathbf{w}_{t+1}),\mathbf{w}-\mathbf{w}_{t+1}\rangle+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{t+1}\|_{H_{t}}^{2}$ satisfies
$L_{t}(\mathbf{w})\leq\tilde{L}_{t}(\mathbf{w})+e^{M^{2}\|\mathbf{w}-\mathbf{w}_{t+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{t+1}\|_{\nabla\ell_{t}(\mathbf{w}_{t+1})}^{2}.$
Lemma F.9.
Let $\mathbf{w}_{t+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\langle\nabla\ell_{t}(\mathbf{w}_{t}),\mathbf{w}\rangle+\frac{1}{2\eta}\|\mathbf{w}-\mathbf{w}_{t}\|_{\tilde{H}_{t}}^{2}$. Then, we have
$\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\|_{H_{t}}\leq 2\eta\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{H_{t}^{-1}}\leq\frac{2\eta}{\lambda}\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq\frac{4\eta}{\lambda}.$
Proof of Lemma F.9.
We make a slight improvement to Lemma 20 of Zhang and Sugiyama [51].
Lemma F.10 (Lemma 20 of Zhang and Sugiyama 51).
Under the same conditions as Lemma F.9, the following holds:
$\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\|_{H_{t}}\leq 2\eta\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{H_{t}^{-1}}\leq\frac{2\eta}{\lambda}\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq\frac{2\eta\sqrt{K}}{\lambda}.$
In the proof of Lemma 20 by Zhang and Sugiyama [51], the bound for $\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}$ was given as $\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq 2\sqrt{K}$. However, a tighter bound can be established: $\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq 2$. Using this tighter bound, we can refine and derive the improved result. ∎
Appendix G Proofs of Theorem 3
In this section, we provide the proof of Theorem 3.
G.1 Adversarial Rewards Construction
We follow a similar context construction as in Section D.1, but now consider the case where all context vectors are distinct and the rewards are non-uniform. Accordingly, the total number of items is $N=\binom{d}{d/4}$. Given $\mathbf{w}_{V}$, we define a unique item $i^{\star}\in[N]$ as the one that maximizes $x_{i}^{\top}\mathbf{w}_{V}$, i.e., $x_{i^{\star}}=x_{V}$. In addition, let $i^{\prime}\in[N]\setminus{i^{\star}}$ denote the item with the second-largest value of $x_{i}^{\top}\mathbf{w}_{V}$. At each round $t$, we assign a reward of $1$ to $i^{\star}$ when $t$ is odd, and a reward of $1$ to one randomly selected item other than $i^{\star}$ when $t$ is even. For simplicity, let ${\mathcal{T}}^{\star}\subseteq[T]$ denote the set of rounds corresponding to the former case, i.e., when the reward $1$ is assigned to $i^{\star}$. Specifically, for $t\in{\mathcal{T}}^{\star}$, the rewards are constructed as follows:
$r_{ti}=\begin{cases}1,&\text{for }i=i^{\star}\\ \nu,&\text{for }i=i^{\prime}\\ \gamma,&\text{otherwise},\end{cases}$
where $\nu$ and $\gamma$ are defined as
$\nu=\frac{e^{d/4-0.5}}{v_{0}+e^{d/4-0.5}},$ (G.1)
and
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}=\frac{\nu}{v_{0}+1}<1.$
Moreover, for rounds $t\notin{\mathcal{T}}^{\star}$, let $\hat{i}\in[N]\setminus{i^{\star}}$ denote the randomly selected item (other than $i^{\star}$) that receives a reward of $1$. The rewards are then defined as follows:
$r_{ti}=\begin{cases}1,&\text{for }i=\hat{i}\\ \nu,&\text{for }i=i^{\prime}\\ \gamma,&\text{otherwise}.\end{cases}$
G.2 Main Proof of Theorem 3
Proof of Theorem 3.
We begin the proof by introducing useful lemmas.
Lemma G.1.
Let $R_{t}(S^{\star}_{t},\mathbf{w}_{V})=\frac{\sum_{i\in S^{\star}_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})r_{ti}}{v_{0}+\sum_{j\in S^{\star}_{t}}\exp(x_{j}^{\top}\mathbf{w}_{V})}$. Then, we have
$S^{\star}_{t}=\begin{cases}\{i^{\star}\}&\text{for }t\in{\mathcal{T}}^{\star}\\ \{\hat{i},i^{\prime}\}&\text{for }t\notin{\mathcal{T}}^{\star}.\end{cases}$
Lemma G.1 establishes that any policy which relies solely on rewards—without estimating the underlying parameter—incurs a regret of order $\Omega(T)$ in this instance. In particular, the reward distributions over rounds in $\mathcal{T}$$^{\star}$ and $T]\setminus{\mathcal{T}}^{\star}$ are identical up to item permutations, making them indistinguishable based solely on the given rewards. Consequently, reward-based heuristics—such as selecting the top-1 or top-2 items, or choosing items whose rewards exceed a fixed threshold—necessarily suffer $\Omega(T)$ regret on $\mathcal{T}$$^{\star}$, $T]\setminus{\mathcal{T}}^{\star}$, or both. For instance, the top-1 strategy incurs zero regret for $t\in{\mathcal{T}}^{\star}$ but $\Omega(T)$ regret for $t\notin{\mathcal{T}}^{\star}$, since the optimal assortment size in that case is 2. Conversely, the top-2 strategy suffers $\Omega(T)$ regret for $t\in{\mathcal{T}}^{\star}$ and zero regret for $t\notin{\mathcal{T}}^{\star}$. Similarly, a threshold-based policy that selects items whose rewards exceed a fixed value yields a fixed-size assortment for all $t$ (since the reward distributions are identical across rounds), and therefore necessarily incurs $\Omega(T)$ regret in one or both cases. Therefore, we can safely exclude such reward-based heuristics from consideration and focus our analysis on the rounds in $\mathcal{T}$$^{\star}$.
Lemma G.2.
For any $t\in{\mathcal{T}}^{\star}$ and $i^{\star}\notin S_{t}$, we have
$R_{t}(S_{t},\mathbf{w}_{V})\leq\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
Lemma G.2 shows that, for $t\in{\mathcal{T}}^{\star}$ and $i^{\star}\notin S_{t}$, the expected revenue of $S_{t}$ can be upper bounded by that of a singleton set whose reward has been replaced with $1$.
Now, we are ready to provide the proof of Theorem 3. Let $x_{U_{1}},\dots,x_{U_{L}}$ be the distinct feature vectors contained in assortments $S_{t}$ with $U_{1},\dots,U_{L}\in\mathcal{V}_{d/4}$. Let $\tilde{U}_{t}$ be the (unique) item among $U_{1},\dots,U_{L}$ that maximizes $x_{U}^{\top}\mathbf{w}_{V}$, i.e., $\tilde{U}_{t}=\mathop{\mathrm{argmax}}_{U\in\{U_{1},\dots,U_{L}\}}x_{U}^{\top}\mathbf{w}_{V}$, where $\mathbf{w}_{V}$ is the underlying parameter. Then, we have
$\sum_{t=1}^{T}$ $R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})$
$\geq\sum_{t\in{\mathcal{T}}^{\star}}R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})$
$=\sum_{t\in{\mathcal{T}}^{\star}}\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-R_{t}(S_{t},\mathbf{w}_{V})$
$\geq\sum_{t\in{\mathcal{T}}^{\star}}\mathbf{1}\{i^{\star}\notin S_{t}\}\left(\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\right)$
$=\sum_{t\in{\mathcal{T}}^{\star}}\mathbf{1}\{i^{\star}\notin S_{t}\}\left(\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-\frac{\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}\right),$
where the first equality holds by Lemma G.1 (and recall that $x_{i^{\star}}=x_{V}$), and the second inequality holds by Lemma G.2. Then, by applying Lemma D.1 with $K=1$, we can further bound the regret as follows:
$\sum_{t=1}^{T}R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})\geq\frac{v_{0}}{(v_{0}+e)^{2}}\sum_{t\in{\mathcal{T}}^{\star}}\frac{\delta\epsilon}{2\sqrt{d}},$
where $\epsilon\in(0,1/d\sqrt{d})$ and $\delta=d/4-|\tilde{U}_{t}\cap V|$. Note that for $i^{\star}\in S_{t}$, we have $\tilde{U}_{t}=V$, and hence we get $\delta=0$.
The remainder of the analysis follows directly from the proof of Theorem 1 with $K=1$ and ${\mathcal{T}}^{\star}=T/2$. Therefore, setting $v_{0}=\Theta(1)$, we obtain
$\sup_{\mathbf{w}}\mathbb{E}_{\mathbf{w}}^{\pi}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]$ $\geq\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})=\Omega\left(d\sqrt{T}\right).$
This concludes the proof of Theorem 3. ∎
G.3 Proofs of Lemmas for Theorem 3
G.3.1 Proof of Lemma G.1
Proof of Lemma G.1.
First, for any $t\in[T]$ and all $i\in S^{\star}_{t}$, we show that $r_{ti}\geq R_{t}(S^{\star}_{t},\mathbf{w}_{V})$. We prove this by contradiction. Assume that there exists $i\in S^{\star}_{t}$ such that $r_{ti}<R_{t}(S^{\star}_{t},\mathbf{w}_{V})$. Then, removing the item $i$ from the assortment $S^{\star}_{t}$ yields higher expected revenue. This contradicts the optimality of $S^{\star}_{t}$. Thus, for any $t\in[T]$, we have
$r_{ti}\geq R_{t}(S^{\star}_{t},\mathbf{w}_{V}),\quad\forall i\in S^{\star}_{t}.$ (G.2)
Hence, for $t\in{\mathcal{T}}^{\star}$, we have $i^{\star}\in S_{t}^{\star}$, since $r_{i^{\star}}=1>R_{t}(S^{\star}_{t},\mathbf{w}_{V})$. Furthermore, for any other item $i\neq i^{\star}$, by definition of $\nu$ in Equation (G.1), we have
$r_{ti}\leq\nu=\frac{e^{d/4-0.5}}{v_{0}+e^{d/4-0.5}}<\frac{e^{d/4}}{v_{0}+e^{d/4}}=\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}=R_{t}(\{i^{\star}\},\mathbf{w}_{V}).$
Therefore, by Equation (G.2), only the item $i^{\star}$ is included in the optimal assortment for all $t\in{\mathcal{T}}^{\star}$.
We next analyze the case $t\notin{\mathcal{T}}^{\star}$. Using the same reasoning as above and applying Equation (G.2), we obtain that $\hat{i}\in S_{t}^{\star}$, since $r_{t\hat{i}}=1>R_{t}(S_{t}^{\star},\mathbf{w}_{V})$. Moreover, we have
$r_{ti^{\prime}}=\nu=\frac{e^{d/4-0.5}}{v_{0}+e^{d/4-0.5}}>\frac{e^{d/4-1}}{v_{0}+e^{d/4-1}}=R_{t}(\{\hat{i}\},\mathbf{w}_{V}),$
which implies $R_{t}(\{\hat{i}\},\mathbf{w}_{V})<R_{t}(\{\hat{i},i^{\prime}\},\mathbf{w}_{V})$. And for any $i\notin\{\hat{i},i^{\prime}\}$,
$r_{ti}=\gamma=\frac{\nu}{v_{0}+1}<\frac{1}{v_{0}+1}<R_{t}(\{\hat{i}\},\mathbf{w}_{V})<R_{t}(\{\hat{i},i^{\prime}\},\mathbf{w}_{V}).$
Therefore, by Equation (G.2), we can conclude that $S^{\star}_{t}=\{\hat{i},i^{\prime}\}$. This completes the proof. ∎
G.3.2 Proof of Lemma G.2
Proof of Lemma G.2.
We provide a proof by considering the following cases:
Case 1. $i^{\prime}\in S_{t}$ and $i^{\star}\notin S_{t}$.
Recall that, by the construction of rewards, we have
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}.$ (G.3)
This implies that
$\left\{\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right\}\frac{\left(v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\right)\nu}{v_{0}+1}\leq\left\{\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right\}\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu$
$\Leftrightarrow\left(\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})+\frac{\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+1}\right)\left(v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\right)$
$\quad\,\,\leq\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\left(v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})+\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right)$
$\Leftrightarrow\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})+\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})/(v_{0}+1)}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}.$ (G.4)
Therefore, we have
$R(S_{t},\mathbf{w}_{V})$ $=\frac{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})r_{ti}}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}=\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu+\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\cdot\nu/(v_{0}+1)}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}$
$\leq\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}\leq\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})},$
where the first inequality holds by Equation (G.4).
Case 2. $i^{\prime},i^{\star}\notin S_{t}$.
Let us return to Equation (G.3). Since $\frac{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\geq 1$, we have
$\gamma\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\cdot\frac{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})},$
which is equivalent to
$\frac{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
Hence, we get
$R(S_{t},\mathbf{w}_{V})$ $=\frac{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}$
$\leq\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
This concludes the proof. ∎
Appendix H Proofs of Theorem 4
In this section, we provide the proof of Theorem 4. Since we now consider the case of non-uniform rewards, the sizes of both the chosen assortment $S_{t}$, and the optimal assortment, $S_{t}^{\star}$ are no longer fixed at $K$.
We begin the proof by introducing additional useful lemmas. Lemma H.1 shows that $\tilde{R}_{t}(S_{t})$, defined in Equation (6), is an upper bound of the true expected revenue of the optimal assortment, $R_{t}(S_{t}^{\star},\mathbf{w}^{\star})$.
Lemma H.1 (Lemma 4 in Oh and Iyengar 39).
Let $\tilde{R}_{t}(S)=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$. And suppose $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$. If for every item $i\in S_{t}^{\star}$, $\alpha_{ti}\geq x_{ti}^{\top}\mathbf{w}^{\star}$, then for all $t\geq 1$, the following inequalities hold:
$R_{t}(S_{t}^{\star},\mathbf{w}^{\star})\leq\tilde{R}_{t}(S_{t}^{\star})\leq\tilde{R}_{t}(S_{t}).$
Note that Lemma H.1 does not claim that the expected revenue is a monotone function in general. Instead, it specifically states that the value of the expected revenue, when associated with the optimal assortment, increases with an increase in the MNL parameters [7, 39].
Lemma H.2 shows that $\tilde{R}_{t}(S_{t})$ increases as the utilities of items in $S_{t}$ increase.
Lemma H.2.
Let $\tilde{R}_{t}(S)=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$ and $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$. Assume $\alpha^{\prime}_{ti}\geq\alpha_{ti}\geq 0$ for all $i\in[N]$. Then, we have
$\tilde{R}_{t}(S_{t})\leq\frac{\sum_{i\in S_{t}}\exp(\alpha^{\prime}_{ti})r_{ti}}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha^{\prime}_{tj})}.$
Furthermore, we provide a novel elliptical potential Lemma H.3 for the centralized context vectors $\tilde{x}_{ti}$.
Lemma H.3.
Let $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$, where $\mathcal{G}_{s}(\mathbf{w})=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w})p_{s}(j|S_{s},\mathbf{w})x_{si}x_{sj}^{\top}$. Define $\tilde{x}_{si}=x_{si}-\mathbb{E}_{j\sim p_{s}(\cdot|S_{s},\mathbf{w}_{s+1})}[x_{sj}]$. Suppose $\lambda\geq 2$. Then the following statements hold true:
  1. (1)
    $\sum_{s=1}^{t}\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})\|\tilde{x}_{si}\|_{H_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{t}{d\lambda}\right)$,
  2. (2)
    $\sum_{s=1}^{t}\max_{i\in S_{s}}\|\tilde{x}_{si}\|_{H_{s}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{t}{d\lambda}\right)$.
Now, we prove the Theorem 4.
H.1 Proof of Theorem 4
Proof of Theorem 4.
Let $\alpha^{\prime}_{ti}=x_{ti}^{\top}\mathbf{w}^{\star}+2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$. If $\mathbf{w}^{\star}\in\mathcal{C}_{t}(\delta)$, then, by Lemma E.1, we have
$\alpha_{ti}\leq x_{ti}^{\top}\mathbf{w}^{\star}+2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}=\alpha^{\prime}_{ti}.$
We denote $\accentset{\approx}{R}_{t}(S_{t})=\frac{\sum_{i\in S_{t}}\exp(\alpha^{\prime}_{ti})r_{ti}}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha^{\prime}_{tj})}$. Then, we can bound the regret as follows:
$\sum_{t=1}^{T}R_{t}(S_{t}^{\star},\mathbf{w}^{\star})-R_{t}(S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\accentset{\approx}{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star}),$
where the first inequality holds by Lemma H.1 and the last inequality holds by Lemma H.2.
Now, we define $\tilde{Q}:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})^{\top}\in\mathbb{R}^{|S_{t}|}$, $\tilde{Q}(\mathbf{u})=\sum_{k=1}^{|S_{t}|}\frac{\exp(u_{k})r_{ti_{k}}}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Let $S_{t}=\{i_{1},\dots,i_{|S_{t}|}\}$. Moreover, for all $t\geq 1$, let $\mathbf{u}_{t}=(u_{ti_{1}},\dots u_{ti_{|S_{t}|}})^{\top}=(\alpha^{\prime}_{ti_{1}},\dots,\alpha^{\prime}_{ti_{{|S_{t}|}}})^{\top}$ and $\mathbf{u}^{\star}_{t}=(u_{ti_{1}}^{\star},\dots u_{ti_{|S_{t}|}}^{\star})^{\top}=(x_{ti_{1}}^{\top}\mathbf{w}^{\star},\dots,x_{ti_{{|S_{t}|}}}^{\top}\mathbf{w}^{\star})^{\top}$. Then, by applying a second order Taylor expansion, we obtain
$\sum_{t=1}^{T}\accentset{\approx}{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})$ $=\sum_{t=1}^{T}\tilde{Q}(\mathbf{u}_{t})-\tilde{Q}(\mathbf{u}^{\star}_{t})$
$=\underbrace{\sum_{t=1}^{T}\nabla\tilde{Q}(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(C)}}+\underbrace{\frac{1}{2}\sum_{t=1}^{T}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})^{\top}\nabla^{2}\tilde{Q}(\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(D)}},$
where $\bar{\mathbf{u}}_{t}=(\bar{u}_{ti_{1}},\dots,\bar{u}_{ti_{|S_{t}|}})^{\top}\in\mathbb{R}^{|S_{t}|}$ is the convex combination of $\mathbf{u}_{t}$ and $\mathbf{u}^{\star}_{t}$.
We first bound the term (C).
$\sum_{t=1}^{T}\nabla\tilde{Q}(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}\frac{\exp(x_{ti}^{\top}\mathbf{w}^{\star})r_{ti}}{v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star})}(u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}\frac{\exp(x_{tj}^{\top}\mathbf{w}^{\star})r_{tj}\sum_{i\in S_{t}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{(v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star}))^{2}}(u_{ti}-u_{ti}^{\star})$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}(u_{ti}-u_{ti}^{\star})-\sum_{i\in S_{t}}\sum_{j\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}p_{t}(j|S_{t},\mathbf{w}^{\star})(u_{tj}-u_{tj}^{\star})$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left((u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})(u_{tj}-u_{tj}^{\star})\right)$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})2\beta_{t}(\delta)\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right).$
Let $x_{t0}=\mathbf{0}$. Then, we can further bound the right-hand side as follows:
$2\sum_{t=1}^{T}\beta_{t}(\delta)$ $\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}\cup\{0\}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\right)$
$\leq 2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\right)$
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\right)$
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\left\|\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[x_{tj}\right]\right\|_{H_{t}^{-1}}\right)$
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|x_{ti}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[x_{tj}\right]\|_{H_{t}^{-1}}$
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left\|x_{ti}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[x_{tj}\right]\right\|_{H_{t}^{-1}},$
where, in the first inequality, we define $S_{t}^{+}\subseteq S_{t}$ as the subset of items in $S_{t}$ such that the term $\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\geq 0$ and $r_{ti}\in[0,1]$, the second inequality holds because $\beta_{1}(\delta)\leq\dots\leq\beta_{T}(\delta)$, the third inequality holds due to Jensen’s inequality, and the second-to-last inequality holds due to the fact that $\|\mathbf{a}\|=\|\mathbf{a}-\mathbf{b}+\mathbf{b}\|\leq\|\mathbf{a}-\mathbf{b}\|+\|\mathbf{b}\|$ for any vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{d}$.
For simplicity, we denote $\mathbb{E}_{\mathbf{w}}[x_{tj}]=\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w})}[x_{tj}]$. Let $\bar{x}_{ti}=x_{ti}-\mathbb{E}_{\mathbf{w}^{\star}}[x_{tj}]$ and $\tilde{x}_{ti}=x_{ti}-\mathbb{E}_{\mathbf{w}_{t+1}}[x_{tj}]$. Then, we have
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|x_{ti}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[x_{tj}\right]\|_{H_{t}^{-1}}=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\bar{x}_{ti}\|_{H_{t}^{-1}}$
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\bar{x}_{ti}-\tilde{x}_{ti}\|_{H_{t}^{-1}}+\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\bar{x}_{ti}-\tilde{x}_{ti}\|_{H_{t}^{-1}}+\sum_{t=1}^{T}\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\mathbf{w}^{\star})-p_{t}(i|S_{t},\mathbf{w}_{t+1})\right)\|\tilde{x}_{ti}\|_{H_{t}^{-1}}$
$+\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}},$ (H.1)
where the inequality holds by the triangle inequality. Now, we bound the terms on the right-hand side of Equation (H.1) individually. For the first term, we have
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\bar{x}_{ti}-\tilde{x}_{ti}\|_{H_{t}^{-1}}=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left\|\mathbb{E}_{\mathbf{w}_{t+1}}[x_{tj}]-\mathbb{E}_{\mathbf{w}^{\star}}[x_{tj}]\right\|_{H_{t}^{-1}}$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left\|\sum_{j\in S_{t}}\left(p_{t}(j|S_{t},\mathbf{w}_{t+1})-p_{t}(j|S_{t},\mathbf{w}^{\star})\right)x_{tj}\right\|_{H_{t}^{-1}},$
where the last equality holds due to the setting of $x_{t0}=\mathbf{0}$. By the mean value theorem, there exists $\bm{\xi}_{t}=(1-c)\mathbf{w}^{\star}+c\mathbf{w}_{t+1}$ for some $c\in(0,1)$ such that
$\left\|\sum_{j\in S_{t}}\left(p_{t}(j|S_{t},\mathbf{w}_{t+1})-p_{t}(j|S_{t},\mathbf{w}^{\star})\right)x_{tj}\right\|_{H_{t}^{-1}}=\left\|\sum_{j\in S_{t}}\nabla p_{t}(j|S_{t},\bm{\xi}_{t})^{\top}(\mathbf{w}_{t+1}-\mathbf{w}^{\star})x_{tj}\right\|_{H_{t}^{-1}}$
$\leq\sum_{j\in S_{t}}\left|\nabla p_{t}(j|S_{t},\bm{\xi}_{t})^{\top}(\mathbf{w}_{t+1}-\mathbf{w}^{\star})\right|\left\|x_{tj}\right\|_{H_{t}^{-1}}$
$=\sum_{j\in S_{t}}\left|\left(p_{t}(j|S_{t},\bm{\xi}_{t})x_{tj}-p_{t}(j|S_{t},\bm{\xi}_{t})\sum_{k\in S_{t}}p_{t}(k|S_{t},\bm{\xi}_{t})x_{tk}\right)^{\top}(\mathbf{w}_{t+1}-\mathbf{w}^{\star})\right|\left\|x_{tj}\right\|_{H_{t}^{-1}}$
$\leq\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left|x_{tj}^{\top}(\mathbf{w}_{t+1}-\mathbf{w}^{\star})\right|\left\|x_{tj}\right\|_{H_{t}^{-1}}$
$+\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}\sum_{k\in S_{t}}p_{t}(k|S_{t},\bm{\xi}_{t})\left|x_{tk}^{\top}(\mathbf{w}_{t+1}-\mathbf{w}^{\star})\right|$
$\leq\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t}}\left\|x_{tj}\right\|_{H_{t}^{-1}}^{2}$
$+\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}\sum_{k\in S_{t}}p_{t}(k|S_{t},\bm{\xi}_{t})\|x_{tk}\|_{H_{t}^{-1}}\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t}}$
$\leq\beta_{t+1}(\delta)\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}^{2}+\beta_{t+1}(\delta)\left(\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}\right)^{2}$
$\leq 2\beta_{t+1}(\delta)\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}^{2}$
$\leq 2\beta_{t+1}(\delta)\max_{j\in S_{t}}\left\|x_{tj}\right\|_{H_{t}^{-1}}^{2},$
where the fourth inequality holds by Lemma 1 and the second-to-last inequality holds due to Jensen’s inequality. Hence, we have
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\bar{x}_{ti}-\tilde{x}_{ti}\|_{H_{t}^{-1}}$ $\leq 2\sum_{t=1}^{T}\beta_{t+1}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\max_{j\in S_{t}}\left\|x_{tj}\right\|_{H_{t}^{-1}}^{2}$
$\leq 2\beta_{T+1}(\delta)\sum_{t=1}^{T}\max_{i\in S_{t}}\left\|x_{ti}\right\|_{H_{t}^{-1}}^{2}$
$\leq\frac{4}{\kappa}\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),$ (H.2)
where the last inequality holds by Lemma E.2. Using similar reasoning, we can bound the second term of Equation (H.1). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
$\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\mathbf{w}^{\star})-p_{t}(i|S_{t},\mathbf{w}_{t+1})\right)\|\tilde{x}_{ti}\|_{H_{t}^{-1}}=\sum_{i\in S_{t}}\nabla p_{t}(i|S_{t},\bm{\xi}^{\prime}_{t})^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}$
$=\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\bm{\xi}^{\prime}_{t})x_{ti}-p_{t}(i|S_{t},\bm{\xi}^{\prime}_{t})\sum_{k\in S_{t}}p_{t}(k|S_{t},\bm{\xi}^{\prime}_{t})x_{tk}\right)^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}$
$\leq\beta_{t+1}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}^{\prime}_{t})\|x_{ti}\|_{H_{t}^{-1}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}$
$+\beta_{t+1}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\bm{\xi}^{\prime}_{t})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}\sum_{k\in S_{t}}p_{t}(k|S_{t},\bm{\xi}^{\prime}_{t})\|x_{tk}\|_{H_{t}^{-1}}$
$\leq\beta_{t+1}(\delta)\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}+\beta_{t+1}(\delta)\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}\max_{k\in S_{t}}\|x_{tk}\|_{H_{t}^{-1}}.$
Then, by applying the AM-GM inequality to each term, we obtain
$\beta_{t+1}(\delta)\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}+\beta_{t+1}(\delta)\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}\max_{k\in S_{t}}\|x_{tk}\|_{H_{t}^{-1}}$
$\leq\beta_{t+1}(\delta)\max_{i\in S_{t}}\frac{\|x_{ti}\|_{H_{t}^{-1}}^{2}+\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}}{2}+\beta_{t+1}(\delta)\frac{\left(\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}\right)^{2}+\left(\max_{k\in S_{t}}\|x_{tk}\|_{H_{t}^{-1}}\right)^{2}}{2}$
$=\beta_{t+1}(\delta)\max_{i\in S_{t}}\frac{\|x_{ti}\|_{H_{t}^{-1}}^{2}+\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}}{2}+\beta_{t+1}(\delta)\frac{\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}+\max_{k\in S_{t}}\|x_{tk}\|_{H_{t}^{-1}}^{2}}{2}$
$\leq 2\beta_{t+1}(\delta)\max\left\{\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\right\}$
where the equality holds since $(\max_{i}a_{i})^{2}=\max_{i}a_{i}^{2}$ for any $a_{i}\geq 0$. Thus, by Lemma H.3 (or Lemma E.2), we get
$\sum_{t=1}^{T}$ $\sum_{i\in S_{t}}\left(p_{t}(i|S_{t},\mathbf{w}^{\star})-p_{t}(i|S_{t},\mathbf{w}_{t+1})\right)\|\tilde{x}_{ti}\|_{H_{t}^{-1}}$
$\leq 2\beta_{t+1}(\delta)\max\left\{\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\right\}\leq\frac{4}{\kappa}\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),,$ (H.3)
Finally, we bound the third term of Equation (H.1). By the Cauchy-Schwarz inequality, we have
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}$ $\leq\sqrt{\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})}\sqrt{\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}}$
$\leq\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)},$ (H.4)
where the last inequality holds by Lemma H.3. Plugging Equation (H.2), (H.3), and (H.4) into Equation (H.1), we get
$\sum_{t=1}^{T}\sum_{i\in S_{t}}$ $p_{t}(i|S_{t},\mathbf{w}^{\star})\|x_{ti}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[x_{tj}\right]\|_{H_{t}^{-1}}$
$\leq\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)}+\frac{8}{\kappa}\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right)$
Thus, we can bound the term (c) as follows:
$\sum_{t=1}^{T}\nabla\tilde{Q}(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})\leq 2\beta_{T}(\delta)\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)}+\frac{16}{\kappa}\beta_{T}(\delta)\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),$ (H.5)
Now, we bound the term (D). Define $Q:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})\in\mathbb{R}^{|S_{t}|}$, $Q(\mathbf{u})=\sum_{i=1}^{|S_{t}|}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Then, we have $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ since $r_{ti}\in[0,1]$. By following the similar reasoning from Equation (E.3) to Equation (E.5) in Section E.1, we have
$\frac{1}{2}\sum_{t=1}^{T}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})^{\top}\nabla^{2}\tilde{Q}(\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})$ $=\frac{1}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t}}(u_{ti}-u_{ti}^{\star})\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}(u_{tj}-u_{tj}^{\star})$
$\leq\frac{1}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t}}|u_{ti}-u_{ti}^{\star}|\left|\frac{\partial^{2}Q}{\partial i\partial j}\right||u_{tj}-u_{tj}^{\star}|$
$\leq 10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}.$ (H.6)
where the first inequality holds because $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$. Combining Equation (H.5) and (H.6), we derive that
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})$ $\leq 2\beta_{T}(\delta)\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)}+\frac{16}{\kappa}\beta_{T}(\delta)\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right)$
$+10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}$
$=\tilde{\mathcal{O}}\left(d\sqrt{T}+\frac{1}{\kappa}d^{2}\right),$
where $\beta_{T}(\delta)=\mathcal{O}\left(\sqrt{d}\log T\log K\right)$. This concludes the proof of Theorem 4. ∎
Remark H.1.
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, since $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right)$ (refer Corollary F.1), we have $\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(B^{3/2}d\sqrt{T}+\frac{1}{\kappa}B^{3}d^{2}\right).$
H.2 Proofs of Lemmas for Theorem 4
H.2.1 Proof of Lemma H.2
Proof of Lemma H.2.
We prove the result by first showing that for any $i\in S_{t}$, we have $r_{ti}\geq\tilde{R}_{t}(S_{t})$. Suppose that there exists $i\in S_{t}$ for which $r_{ti}<\tilde{R}_{t}(S_{t})$. Removing item $i$ from the assortment $S_{t}$ results in a higher expected revenue. Consequently, $S_{t}\neq\mathop{\mathrm{argmax}}_{S\in{\mathcal{S}}}\tilde{R}_{t}(S)$, which contradicts the optimality of $S_{t}$. Thus, we have
$r_{ti}\geq\tilde{R}_{t}(S_{t}),\quad\forall i\in S_{t}.$
If we increase $\alpha_{ti}$ to $\alpha_{ti}^{\prime}$ for all $i\in S_{t}$, the probability of selecting the outside option decreases. In other words, the sum of probabilities of choosing any $i\in S_{t}$ increases. Since $r_{ti}\geq\tilde{R}_{t}(S_{t})$ for all $i\in S_{t}$, this results in an increase in revenue. Hence, we get
$\tilde{R}_{t}(S_{t})=\frac{\sum_{i\in S_{t}}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha_{tj})}\leq\frac{\sum_{i\in S_{t}}\exp(\alpha^{\prime}_{ti})r_{ti}}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha^{\prime}_{tj})}.$
This concludes the proof. ∎
H.2.2 Proof of Lemma H.3
Proof of Lemma H.3.
For notational simplicity, let $\mathbb{E}_{\mathbf{w}}[x_{tj}]=\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w})}[x_{tj}]$. Let $x_{t0}=\mathbf{0}$. We can rewrite $\mathcal{G}_{s}(\mathbf{w})$ in the following way:
$\mathcal{G}_{s}(\mathbf{w}_{s+1})$
$=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(j|S_{s},\mathbf{w}_{s+1})x_{si}x_{sj}^{\top}$ (H.7)
$=\sum_{i\in S_{s}\cup\{0\}}p_{s}(i|S_{s},\mathbf{w}_{s+1})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}\cup\{0\}}\sum_{j\in S_{s}\cup\{0\}}p_{s}(i|S_{s},\mathbf{w}_{s+1})p_{s}(j|S_{s},\mathbf{w}_{s+1})x_{si}x_{sj}^{\top}$
$=\mathbb{E}_{\mathbf{w}_{s+1}}[x_{si}x_{si}^{\top}]-\mathbb{E}_{\mathbf{w}_{s+1}}[x_{si}]\left(\mathbb{E}_{\mathbf{w}_{s+1}}[x_{si}]\right)^{\top}$
$=\mathbb{E}_{\mathbf{w}_{s+1}}\left[(x_{si}-\mathbb{E}_{\mathbf{w}_{s+1}}[x_{sm}])(x_{si}-\mathbb{E}_{\mathbf{w}_{s+1}}[x_{sm}])^{\top}\right]$
$=\mathbb{E}_{\mathbf{w}_{s+1}}[\tilde{x}_{si}\tilde{x}_{si}^{\top}]=\sum_{i\in S_{s}\cup\{0\}}p_{s}(i|S_{s},\mathbf{w}_{s+1})\tilde{x}_{si}\tilde{x}_{si}^{\top}\succeq\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})\tilde{x}_{si}\tilde{x}_{si}^{\top}.$
This means that
$H_{t+1}=H_{t}+\mathcal{G}_{t}(\mathbf{w}_{t+1})\succeq H_{t}+\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\tilde{x}_{ti}\tilde{x}_{ti}^{\top}.$ (H.8)
Hence, we can derive that
$\det\left(H_{t+1}\right)\geq\det\left(H_{t}\right)\left(1+\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\right).$
Since $\lambda\geq 1$, for all $t\geq 1$ we have $\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\leq\frac{1}{\lambda}\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{2}\leq 1$. Then, using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
$\sum_{s=1}^{t}\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}_{s+1})\|\tilde{x}_{si}\|_{H_{s}^{-1}}^{2}$ $\leq 2\sum_{s=1}^{t}\log\left(1+p_{s}(i|S_{s},\mathbf{w}_{s+1})\|\tilde{x}_{si}\|_{H_{s}^{-1}}^{2}\right)$
$\leq 2\sum_{s=1}^{t}\log\left(\frac{\det(H_{s+1})}{\det(H_{s})}\right)$
$=2\log\left(\frac{\det(H_{t+1})}{\det(H_{1})}\right)$
$\leq 2d\log\left(\frac{\mathop{\mathrm{tr}}(H_{t+1})}{d\lambda}\right)\leq 2d\log\left(1+\frac{t}{d\lambda}\right).$
This proves the first inequality.
To show the second inequality, we come back to Equation (H.8). By the definition of $\kappa$, we get
$H_{t+1}$ $=H_{t}+\mathcal{G}_{t}(\mathbf{w}_{t+1})=H_{t}+\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\tilde{x}_{ti}\tilde{x}_{ti}^{\top}$
$\succeq H_{t}+\kappa\sum_{i\in S_{t}}\tilde{x}_{ti}\tilde{x}_{ti}^{\top}.$
Thus, we obtain that
$\det\left(H_{t+1}\right)\geq\det\left(H_{t}\right)\left(1+\kappa\sum_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\right)\geq\det\left(H_{t}\right)\left(1+\kappa\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\right).$
Since $\lambda\geq 1$, for all $t\geq 1$ we have $\kappa\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\leq\frac{\kappa}{\lambda}\|\tilde{x}_{ti}\|_{2}\leq\kappa$. We then reach the conclusion in the same manner:
$\sum_{s=1}^{t}\max_{i\in S_{s}}\|\tilde{x}_{si}\|_{H_{s}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{s=1}^{t}\log\left(1+\kappa\max_{i\in S_{s}}\|\tilde{x}_{si}\|_{H_{s}^{-1}}^{2}\right)$
$\leq\frac{2}{\kappa}\sum_{s=1}^{t}\log\left(\frac{\det(H_{s+1})}{\det(H_{s})}\right)$
$=\frac{2}{\kappa}\log\left(\frac{\det(H_{t+1})}{\det(H_{1})}\right)$
$\leq\frac{2}{\kappa}d\log\left(\frac{\mathop{\mathrm{tr}}(H_{t+1})}{d\lambda}\right)\leq\frac{2}{\kappa}d\log\left(1+\frac{t}{d\lambda}\right).$
This proves the second inequality. ∎
Appendix I Proof of Proposition 1
I.1 Main Proof of Proposition 1
Proof of Proposition 1.
We construct the parameters and features in the same way as in Section D.1. Recall that, by Proposition D.1, it is sufficient to bound $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$. We denote $\tilde{U}_{t}$ as the unique $U^{\star}\in\mathcal{V}_{d/4}$ in $\tilde{S}_{t}$. We define a function $\mu:\mathbb{R}\rightarrow[0,1]$ such that for any $z\in\mathbb{R}$, $\mu(z)=\frac{K\exp(z)}{v_{0}+K\exp(z)}$. Since all items in $S^{\star}$ (and in $\tilde{S}_{t}$) are identical, we can express $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})$ (and $\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$) as follows:
$\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})$ $=\frac{K\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+K\exp(x_{V}^{\top}\mathbf{w}_{V})}=\mu(x_{V}^{\top}\mathbf{w}_{V}),$
$\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ $=\frac{K\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}{v_{0}+K\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}=\mu(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V}).$
Given $\mathbf{w}_{V}$, we define $\kappa^{\star}_{t}(\mathbf{w}_{V}):=\sum_{i\in S^{\star}}p_{t}(i|S^{\star},\mathbf{w}_{V})p_{t}(0|S^{\star},\mathbf{w}_{V})$. Since the context vectors $\{x_{ti}\}$ are constructed to be invariant across rounds $t$, $\kappa^{\star}_{t}(\mathbf{w}_{V})$ is also independent of $t$, i.e., $\kappa^{\star}_{1}(\mathbf{w}_{V})=\dots=\kappa^{\star}_{T}(\mathbf{w}_{V})$. Therefore, we omit the index $t$ for simplicity. Note that, in our instance construction, $\kappa^{\star}(\mathbf{w}_{V})=\mu(x_{V}^{\top}\mathbf{w}_{V})\left(1-\mu(x_{V}^{\top}\mathbf{w}_{V})\right)=\dot{\mu}(x_{V}^{\top}\mathbf{w}_{V})$.
We hypothesize that for all $V\in\mathcal{V}_{d/4}$, the regret is dominated by $d\sqrt{\kappa^{\star}(\mathbf{w}_{V})T}$. If this assumption does not hold, then by definition, there exists some $V\in\mathcal{V}_{d/4}$ such that $\mathbb{E}_{V}\left[\textup{{Reg}}_{T}(\mathbf{w}_{V})\right]=\Omega\left(d\sqrt{\kappa^{\star}(\mathbf{w}_{V})T}\right)$, thereby completing the proof.
Hypothesis. There exists a constant $C>0$ such that:
$\mathbb{E}_{V}\left[\textup{{Reg}}_{T}(\mathbf{w}_{V})\right]\leq C\cdot d\sqrt{\kappa^{\star}(\mathbf{w}_{V})T},\quad\,\forall V\in\mathcal{V}_{d/4}.$ (I.1)
Additionally, we set $\mathbf{w}^{\star}=\mathop{\mathrm{argmax}}_{\mathbf{w}_{V}}\kappa^{\star}(\mathbf{w}_{V})$, thus, we have $\kappa^{\star}=\kappa^{\star}(\mathbf{w}^{\star})=\max_{\mathbf{w}_{V}}\kappa^{\star}(\mathbf{w}_{V})$.
To establish an instance-dependent lower bound for $\mu(x_{V}^{\top}\mathbf{w}_{V})-\mu(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})$, we use the following lemma in place of Lemma D.1:
Lemma I.1.
Suppose $\epsilon\in(0,1/d\sqrt{d})$ and define $\delta:=d/4-|\tilde{U}_{t}\cap V|$. Then, we have
$\mu(x_{V}^{\top}\mathbf{w}_{V})-\mu(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\delta\epsilon}{\sqrt{d}}.$
For any $j\in V$, define random variables $\tilde{M}_{j}:=\sum_{t=1}^{T}\mathbf{1}\{j\in\tilde{U}_{t}\}$. Then, by Lemma I.1, for all $V\in\mathcal{V}_{d/4}$, we get
$\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\epsilon}{\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (I.2)
Additionally, we define $\mathcal{V}_{d/4}^{(j)}:=\{V\in\mathcal{V}_{d/4}:j\in V\}$ and $\mathcal{V}_{d/4-1}:=\{V\subseteq[d]:|V|=d/4-1\}$. By averaging both sides of Equation (I.2) with respect to all $V\in\mathcal{V}_{d/4}$, and by following reasoning similar to that in the proof of Theorem 1, we get
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
$\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\epsilon}{\sqrt{d}}\left(\frac{dT}{6}-\max_{V\in\mathcal{V}_{d/4-1}}\sum_{j\notin V}\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|\right).$ (I.3)
For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$. Then, we can bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in Equation (I.3) for any $V\in\mathcal{V}_{d/4-1}$.
$\left|\mathbb{E}_{P}[\tilde{M}_{j}]-\mathbb{E}_{Q}[\tilde{M}_{j}\right]|$ $\leq\sum_{t=0}^{T}t\cdot\left|P[\tilde{M}_{j}=t]-Q[\tilde{M}_{j}=t]\right|$
$\leq T\cdot\|P-Q\|_{\operatorname{TV}}\leq T\cdot\sqrt{\frac{1}{2}\operatorname{KL}(P\|Q)},$ (I.4)
where $\|P-Q\|_{\operatorname{TV}}=\sup_{A}|P(A)-Q(A)|$ | is the total variation distance between $P$ and $Q$, $\operatorname{KL}(P\|Q)=\int(\log\mathrm{d}P/\mathrm{d}Q)\mathrm{d}P$ is s the Kullback-Leibler (KL) divergence between $P$ and $Q$, and the last inequality holds by Pinsker’s inequality. We can derive the instance-dependent bound for the KL divergence term in Equation (I.4) by the following lemma:
Lemma I.2.
For any $V\in\mathcal{V}_{d/4-1}$ and $j\in[d]$, there exists a positive constant $C_{\operatorname{KL}}>0$ such that
$\sum_{j=1}^{d}\operatorname{KL}(P_{V}\|Q_{V\cup\{j\}})\leq C_{\operatorname{KL}}\cdot\epsilon^{2}\kappa^{\star}(\mathbf{w}_{V})T.$
Plugging Equation (I.4) into Equation (I.3), we get
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
$\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\epsilon}{\sqrt{d}}\left(\frac{dT}{6}-T\sum_{j=1}^{d}\sqrt{\frac{1}{2}\operatorname{KL}(P\|Q)}\right)$
$\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\epsilon}{\sqrt{d}}\left(\frac{dT}{6}-T\sqrt{d}\cdot\sqrt{\frac{1}{2}\sum_{j=1}^{d}\operatorname{KL}(P\|Q)}\right)$
$\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\epsilon}{\sqrt{d}}\left(\frac{dT}{6}-T\sqrt{d}\cdot\sqrt{\frac{1}{2}C_{\operatorname{KL}}\cdot\epsilon^{2}\kappa^{\star}(\mathbf{w}_{V})T}\right),$
where the second inequality is due to the Cauchy-Schwartz inequality, and the last inequality is by Lemma I.2.
By setting $\epsilon=\sqrt{\frac{d}{72C_{\operatorname{KL}}\cdot\kappa^{\star}(\mathbf{w}_{V})T}}$, we have
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
$\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{1}{\sqrt{d}}\cdot\sqrt{\frac{d}{72C_{\operatorname{KL}}\cdot\kappa^{\star}(\mathbf{w}_{V})T}}\cdot\frac{dT}{12}$
$=\Omega\left(d\sqrt{\kappa^{\star}(\mathbf{w}_{V})T}\right).$
Recall that by construction, $\kappa^{\star}=\max_{\mathbf{w}_{V}}\kappa^{\star}(\mathbf{w}_{V})$. Thus, by taking the maximum over $\mathbf{w}_{V}$, we complete the proof of Proposition 1. ∎
I.2 Proofs of Lemmas for Proposition 1
I.2.1 Proof of Lemma I.1
Proof of Lemma I.1.
For simplicity, let $x=x_{V}$ and $\hat{x}=x_{\tilde{U}_{t}}$. Then, by the mean value theorem, we have
$\mu(x^{\top}\mathbf{w}_{V})-\mu(\hat{x}^{\top}\mathbf{w}_{V})$ $=\int_{v=0}^{1}\dot{\mu}\left(x^{\top}\mathbf{w}_{V}+v(\hat{x}-x)^{\top}\mathbf{w}_{V}\right)\left(x-\hat{x}\right)^{\top}\mathbf{w}_{V}$
$\geq\frac{\dot{\mu}(x^{\top}\mathbf{w}_{V})}{1+\left|(x-\hat{x})^{\top}\mathbf{w}_{V}\right|}\left(x-\hat{x}\right)^{\top}\mathbf{w}_{V}$
$\geq\frac{\dot{\mu}(x^{\top}\mathbf{w}_{V})}{3}\left(x-\hat{x}\right)^{\top}\mathbf{w}_{V}$
$=\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\left(x-\hat{x}\right)^{\top}\mathbf{w}_{V}$
$\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\delta\epsilon}{\sqrt{d}},$
where the first inequality holds by Lemma I.3, the second inequality holds by the bounded assumption (Assumption 1), and the last inequality holds by the definition of $\delta=d/4-|\tilde{U}_{t}\cap V|$. ∎
I.2.2 Proof of Lemma I.2
Proof of Lemma I.2.
Consider a fixed round $t$, an assortment $\tilde{S}_{t}$, and the set $\tilde{U}_{t}$. Let $U=\tilde{U}_{t}$. Define $m_{j}(\tilde{S}_{t}):=\sum_{x_{U}\in\tilde{S}_{t}}\mathbf{1}\{j\in U\}/K$, which captures the average presence of $j\in U$ across the assortment $\tilde{S}_{t}$. For simplicity, we denote $p=\frac{K\exp(x_{U}^{\top}\mathbf{w}_{V})}{v_{0}+K\exp(x_{U}^{\top}\mathbf{w}_{V})}$ and $q=\frac{K\exp(x_{U}^{\top}\mathbf{w}_{V\cup\{j\}})}{v_{0}+K\exp(x_{U}^{\top}\mathbf{w}_{V\cup\{j\}})}$. Then, we get
$\operatorname{KL}\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)$ $\leq\chi^{2}\!\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)$
$=\chi^{2}\!\left(\operatorname{Bernoulli}\left(p\right)\|\operatorname{Bernoulli}\left(q\right)\right)$
$=\frac{(p-q)^{2}}{q}+\frac{(p-q)^{2}}{1-q}$
$=\frac{(p-q)^{2}}{q(1-q)}$
$=\frac{\left(\mu(x_{U}^{\top}\mathbf{w}_{V})-\mu(x_{U}^{\top}\mathbf{w}_{V\cup\{j\}})\right)^{2}}{\dot{\mu}(x_{U}^{\top}\mathbf{w}_{V\cup\{j\}})}$
$=\frac{\left(\dot{\mu}(x_{U}^{\top}\bar{\mathbf{w}}_{V,j})\right)^{2}}{\dot{\mu}(x_{U}^{\top}\mathbf{w}_{V\cup\{j\}})}\left(x_{U}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right)^{2},$
where in the first inequality, we used $\operatorname{KL}\leq\chi^{2}$ (Tsybakov [48], Chapter 2), where $\chi^{2}$ is a chi-square divergence, the second equality holds by the expression of the chi-square divergence for Bernoulli random variables, and the last equality holds by the mean value theorem, where $\bar{\mathbf{w}}_{V,j}$ is a convex combination of $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup\{j\}}$. Then, by Lemma I.4, we can further bound the last term as follows:
$\operatorname{KL}\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)$ $\leq\dot{\mu}(x_{U}^{\top}\mathbf{w}_{V})e^{3\left|x_{U}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right|}\left(x_{U}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right)^{2}$
$\leq\dot{\mu}(x_{U}^{\top}\mathbf{w}_{V})e^{3/d^{2}}\left(x_{U}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right)^{2}$
$\leq\dot{\mu}(x_{U}^{\top}\mathbf{w}_{V})e^{1/4}\left(x_{U}^{\top}\left(\mathbf{w}_{V}-\mathbf{w}_{V\cup\{j\}}\right)\right)^{2}$
$\leq\dot{\mu}(x_{U}^{\top}\mathbf{w}_{V})e^{1/4}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$
where the second-to-the last inequality holds based on the assumption that $d\geq 4$, and the last inequality holds since $m_{j}(\tilde{S}_{t})\leq 1$.
Now, let us consider $t$ varying over the round $t\in[T]$. Then, we obtain
$\sum_{j=1}^{d}\operatorname{KL}(P_{V}\|Q_{V\cup\{j\}})$ $=\sum_{j=1}^{d}\sum_{t=1}^{T}\mathbb{E}_{V}\left[\operatorname{KL}\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)\right]$
$\leq e^{1/4}\cdot\frac{\epsilon^{2}}{d}\sum_{j=1}^{d}\sum_{t=1}^{T}\mathbb{E}_{V}\left[\dot{\mu}(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})\cdot m_{j}(\tilde{S}_{t})\right]$
$=e^{1/4}\cdot\frac{\epsilon^{2}}{d}\mathbb{E}_{V}\left[\sum_{t=1}^{T}\left(\dot{\mu}(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})\sum_{j=1}^{d}m_{j}(\tilde{S}_{t})\right)\right]$
$\leq e^{1/4}\cdot\frac{\epsilon^{2}}{4}\mathbb{E}_{V}\left[\sum_{t=1}^{T}\dot{\mu}(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})\right]$
$\leq e^{1/4}\cdot\frac{\epsilon^{2}}{4}\left(\dot{\mu}(x_{V}^{\top}\mathbf{w}_{V})T+\mathbb{E}_{V}\left[\textup{{Reg}}_{T}(\mathbf{w}_{V})\right]\right)$
$=e^{1/4}\cdot\frac{\epsilon^{2}}{4}\left(\kappa^{\star}(\mathbf{w}_{V})T+\mathbb{E}_{V}\left[\textup{{Reg}}_{T}(\mathbf{w}_{V})\right]\right)$
$\leq C_{\operatorname{KL}}\cdot\frac{\epsilon^{2}}{2}\left(\kappa^{\star}(\mathbf{w}_{V})T+d\sqrt{\kappa^{\star}(\mathbf{w}_{V})T}\right),$
the equality follows from the chain rule of relative entropy (cf. Exercise 14.11 of Lattimore and Szepesvári [30]), the second inequality holds because $\sum_{j=1}^{d}m_{j}(\tilde{S}_{t})\leq\frac{d}{4}$, second-to-the last inequality holds by Lemma I.5, and the last inequality holds by the hypothesis (Equation (I.1)) with $C_{\operatorname{KL}}>0$.
Furthermore, by the definition of $\kappa$ (Assumption 2), we know that $\kappa^{\star}(\mathbf{w}_{V})\geq\kappa$. Hence, given the assumption that $T\geq d^{2}/\kappa\geq d^{2}/\kappa^{\star}(\mathbf{w}_{V})$, we derive that
$\sum_{j=1}^{d}\operatorname{KL}(P_{V}\|Q_{V\cup\{j\}})$ $\leq C_{\operatorname{KL}}\cdot\epsilon^{2}\kappa^{\star}(\mathbf{w}_{V})T,$
which concludes the proof. ∎
I.3 Technical Lemmas for Proposition 1
Lemma I.3 (Lemma 7 of Abeille et al. [3]).
Let $f$ be a strictly increasing function such that $|\ddot{f}|\leq\dot{f}$, and let $\mathcal{Z}$ be any bounded interval of $\mathbb{R}$. Then, for all $z_{1},z_{2}\in\mathcal{Z}$:
$\int_{v=0}^{1}\dot{f}\left(z_{1}+v(z_{2}-z_{1})\right)dv\geq\frac{\dot{f}(z)}{1+|z_{1}-z_{2}|},\quad\text{for}\,\,z\in\{z_{1},z_{2}\}.$
Lemma I.4 (Lemma 9 of Abeille et al. [3]).
Let $f$ be a strictly increasing function such that $|\ddot{f}|\leq\dot{f}$, and let $\mathcal{Z}$ be any bounded interval of $\mathbb{R}$. Then, for all $z_{1},z_{2}\in\mathcal{Z}$:
$\dot{f}(z_{2})\exp(-|z_{2}-z_{1}|)\leq\dot{f}(z_{1})\leq\dot{f}(z_{2})\exp(|z_{2}-z_{1}|).$
Lemma I.5 (Lemma 11 of Perivier and Goyal [42]).
Let $\kappa_{t}^{\star}:=\sum_{i\in S_{t}^{\star}}p_{t}(i|S_{t}^{\star},\mathbf{w}^{\star})p_{t}(0|S_{t}^{\star},\mathbf{w}^{\star})$. Then, we have
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}(\mathbf{w}^{\star}).$
Appendix J Proof of Proposition 2
In this section, we provide the proof of Proposition 2. We introduce useful lemmas to support the proof.
Lemma J.1 (Theorem 4 of Tran-Dinh et al. [47]).
Let $f:\mathbb{R}^{K}\rightarrow\mathbb{R}$ be a $M_{f}$-self-concordant-like function and let $x,y\in\operatorname{dom}(f)$, then:
$e^{-M_{f}\|y-x\|_{2}}\nabla^{2}f(x)\preceq\nabla^{2}f(y).$
Proof of Proposition 2.
We begin the proof with Equation (E.2) from the proof of Theorem 2:
$\sum_{t=1}^{T}\nabla Q(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}_{t}^{\star})\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}.$ (J.1)
We then divide the total rounds into two disjoint sets, $J_{1}$ and $J_{2}$, such that $J_{1}\bigcup J_{2}=[T]$. Specifically, let $J_{1}=\{t\in[T]|\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\geq\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\}$ and $J_{2}=[T]\setminus J_{1}$. For a better presentation, we also define that:
$g_{t}(S;\mathbf{w})$ $:=\sum_{i\in S}p_{t}(i|S,\mathbf{w})p_{t}(0|S,\mathbf{w})\|x_{ti}\|_{H_{t}^{-1}}.$
Then, we can rewrite the right-hand side of Equation (J.1) as follows:
$2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}$ $p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}$
$=2\beta_{T}(\delta)\sum_{t=1}^{T}g_{t}(S_{t};\mathbf{w}^{\star})$
$=2\beta_{T}(\delta)\sum_{t\in J_{1}}g_{t}(S_{t};\mathbf{w}^{\star})+2\beta_{T}(\delta)\sum_{t\in J_{2}}g_{t}(S_{t};\mathbf{w}^{\star}).$
To bound $\sum_{t\in J_{1}}g_{t}(S_{t};\mathbf{w}^{\star})$, by the mean value theorem, we get
$\sum_{i\in J_{1}}g_{t}(S_{t};\mathbf{w}^{\star})=\sum_{i\in J_{1}}g_{t}(S_{t};\mathbf{w}_{t+1})+\sum_{i\in J_{1}}\nabla_{\mathbf{w}}g_{t}(S_{t};\bar{\mathbf{w}}_{t})^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1}),$ (J.2)
where $\bar{\mathbf{w}}_{t}$ is the convex combination of $\mathbf{w}^{\star}$ and $\mathbf{w}_{t+1}$. The first term of Equation (J.2) can be bound by
$\sum_{t\in J_{1}}g_{t}(S_{t};\mathbf{w}_{t+1})$
$\leq\sqrt{\sum_{t\in J_{1}}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})}\sqrt{\sum_{t\in J_{1}}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})p_{t}(0|S_{t},\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}^{2}}$
$\leq\sqrt{\sum_{t\in J_{1}}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})}\cdot\tilde{\mathcal{O}}\left(\sqrt{d}\right)\leq\sqrt{\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}(\mathbf{w}^{\star})}\cdot\tilde{\mathcal{O}}\left(\sqrt{d}\right),$ (J.3)
where the second-to-the last inequality holds by the definition of $J_{1}$ and the elliptical potential lemma (Lemma E.2), and the last inequality holds by the following Lemma:
Moreover, the second term of Equation (J.2) can be bounded as follows:
$\left|\sum_{i\in J_{1}}\nabla_{\mathbf{w}}g_{t}(S_{t};\bar{\mathbf{w}}_{t})^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\right|$
$=\Bigg|\sum_{t\in J_{1}}p_{t}(0|S_{t},\bar{\mathbf{w}}_{t})\sum_{i\in S_{t}}p_{t}(i|S_{t},\bar{\mathbf{w}}_{t})x_{ti}^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\|x_{ti}\|_{H_{t}^{-1}}$
$-2p_{t}(0|S_{t},\bar{\mathbf{w}}_{t})\sum_{i\in S_{t}}p_{t}(i|S_{t},\bar{\mathbf{w}}_{t})\|x_{ti}\|_{H_{t}^{-1}}\sum_{j\in S_{t}}p_{t}(j|S_{t},\bar{\mathbf{w}}_{t})x_{tj}^{\top}(\mathbf{w}^{\star}-\mathbf{w}_{t+1})\Bigg|$
$\leq\beta_{T}(\delta)\sum_{t\in J_{1}}\sum_{i\in S_{t}}p_{t}(i|S_{t},\bar{\mathbf{w}}_{t})\|x_{ti}\|_{H_{t}^{-1}}^{2}+2\beta_{T}(\delta)\sum_{t\in J_{1}}\left(\sum_{i\in S_{t}}p_{t}(i|S_{t},\bar{\mathbf{w}}_{t})\|x_{ti}\|_{H_{t}^{-1}}\right)^{2}$
$\leq\beta_{T}(\delta)\sum_{t\in J_{1}}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}+2\beta_{T}(\delta)\sum_{t\in J_{1}}\left(\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}\right)^{2}$
$\leq 3\beta_{T}(\delta)\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}=\tilde{\mathcal{O}}\left(d^{3/2}/\kappa\right),$ (J.4)
where the first inequality holds by Lemma 1, and the last inequality holds by Lemma E.2. Combining Equation (J.3) and (J.4), we can bound the term $\sum_{t\in J_{1}}g_{t}(S_{t};\mathbf{w}^{\star})$ as follows:
$\sum_{t\in J_{1}}g_{t}(S_{t};\mathbf{w}^{\star})\leq\tilde{\mathcal{O}}\left(\sqrt{d}\cdot\sqrt{\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}(\mathbf{w}^{\star})}+d^{3/2}/\kappa\right).$ (J.5)
Now, we bound $\sum_{t\in J_{2}}g_{t}(S_{t};\mathbf{w}^{\star})$. We define $H_{t}^{\star}:=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}^{\star})$. Recall that $\nabla^{2}\ell_{s}(\mathbf{w})=\mathcal{G}_{s}(\mathbf{w})$ and $\ell_{s}$ is $3\sqrt{2}$-self-concordant-like function (Proposition C.1). Then, by Lemma J.1, we get $H_{t}\succeq H_{t}^{\star}e^{-3\sqrt{2}}$. With this fact, we can now bound the term $\sum_{t\in J_{2}}g_{t}(S_{t};\mathbf{w}^{\star})$:
$\sum_{t\in J_{2}}g_{t}(S_{t};\mathbf{w}^{\star})$ $\leq\sqrt{\sum_{t\in J_{2}}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})}\sqrt{\sum_{t\in J_{2}}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}^{2}}$
$\leq\sqrt{\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}}\sqrt{\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}^{2}}$
$\leq\sqrt{\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}}\sqrt{e^{3\sqrt{2}}\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{\left(H_{t}^{\star}\right)^{-1}}^{2}}$
$\leq\tilde{\mathcal{O}}\left(\sqrt{d}\cdot\sqrt{\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}}\right),$ (J.6)
where the second inequality follows from Lemma I.5, second-to-the last inequality holds because $H_{t}\succeq H_{t}^{\star}e^{-3\sqrt{2}}$, and the last inequality is obtained by a slight modification of Lemma E.2 with $\mathbf{w}_{1},\dots,\mathbf{w}_{T}=\mathbf{w}^{\star}$.
Now, by plugging Equation (J.5) and (J.6) into Equation (J.1), and using the fact that $\beta_{T}(\delta)=\tilde{\mathcal{O}}(\sqrt{d})$, we obtain that
$2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}}=\tilde{\mathcal{O}}\left(d\sqrt{\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}(\mathbf{w}^{\star})}+d^{3/2}/\kappa\right).$
Furthermore, by applying the same analysis as in the proof of Theorem 2 in order to bound the second-order term in the Taylor expansion (term (B) in Equation (E.1)), we get
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(d\sqrt{\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}(\mathbf{w}^{\star})}+d^{2}/\kappa\right).$
By solving the above equation, we conclude the proof of Proposition 2. ∎
Appendix K Experiment Details and Additional Results
Figure K.1: Runtime per round under uniform rewards (first row) and non-uniform rewards (second row).
Figure K.2: Cumulative regret under uniform rewards with $v_{0}=\Theta(K)$.
For each instance, we sample the true parameter $\mathbf{w}^{\star}$ from a uniform distribution in $1/\sqrt{d},1/\sqrt{d}]^{d}$. For the context features $x_{ti}$, we sample each $x_{ti}$ independently and identically distributed (i.i.d.) from a multivariate Gaussian distribution $\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d})$ and clip it to range $1/\sqrt{d},1/\sqrt{d}]^{d}$. Therefore, we ensure that $\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|x_{ti}\|_{2}\leq 1$, satisfying Assumption 1. For each experimental configuration, we conducted $20$ independent runs for each instance and reported the average cumulative regret (Figure 1) and runtime per round (Figure K.1) for each algorithm. The error bars in Figure 1 and K.2 represent the standard deviations (1-sigma error). We have omitted the error bars in Figure K.1 because they are minimal.
In the uniform reward setting where $r_{ti}=1$, the combinatorial optimization step to select the assortment simply involves sorting items by their utility estimate. In contrast, in the non-uniform reward setting, rewards are sampled from a uniform distribution in each round, i.e., $r_{ti}\sim\operatorname{Unif}(0,1)$. For combinatorial optimization in this setting, we solve an equivalent linear programming (LP) problem that is solvable in polynomial-time [45, 19]. The experiments are run on Xeon(R) Gold 6226R CPU @ 2.90GHz (16 cores).
Figure K.1 presents additional empirical results on the runtime per round. Our algorithm OFU-MNL+ demonstrates a constant computation cost for each round, while the other algorithms exhibit a linear dependence on $t$. It is also noteworthy that the runtime for uniform rewards is approximately $10$ times faster than that for non-uniform rewards. This difference arises because we use linear programming (LP) optimization for assortment selection in the non-uniform reward setting, which is more computationally intensive.
Furthermore, Figure K.2 illustrates the cumulative regrets of the proposed algorithm compared to other baseline algorithms under uniform rewards with $v_{0}=K/5$. Since $v_{0}$ is proportional to $K$, an increase in $K$ does not improve the regret. This observation is also consistent with our theoretical results.
Appendix L Technical Errors in Agrawal et al. [4]
In this section, we discuss the technical errors in Agrawal et al. [4]. There are two main significant errors: There are mainly two significant errors:
1. Equation (16).
$\alpha_{i}(\mathbf{X}_{S_{t}},\theta_{t},\theta^{\star})x_{ti}:=\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta^{\star})-\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta_{t}),$ (L.1)
where $\mathbf{X}_{S_{t}}$ is a design matrix whose columns are the attribute vectors $x_{ti}$ of the items in the assortment $S_{t}$ and $\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta)=P_{t}(i|S_{t},\theta)$.
It appears that the authors may have intended to derive this equation using a first-order exact Taylor expansion. However, in MNL bandits, this equation generally does not hold. Consider a counterexample where $x_{ti}=0$, $x_{tj}\neq 0$ for $j\neq i$, and $\theta^{\star}\neq\theta_{t}$. Then, the left-hand side of Equation (L.1) is equals to 0 (since $x_{ti}=0$), but the right-hand side of Equation (L.1) is not 0, because the denominators of each $\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta^{\star})$ and $\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta_{t})$ differ. This equation only holds in special cases, such as when $K=1$, which corresponds to the logistic bandit case. Equation (16) in Agrawal et al. [4] serves as the foundation for the entire proof in their paper. Consequently, all subsequent results derived from it are also incorrect.
2. Cauchy-Schwarz inequality in the regret analysis (Page 46)
When using the Cauchy-Schwarz inequality on the regret before applying the elliptical potential lemma, they indeed incur an additional $\sqrt{K}$ factor:
$\sum_{t=1}^{T}\min\left(\sum_{i\in S_{t}}\|x_{ti}\|_{\mathbf{J}_{t}^{-1}},1\right)\leq\sqrt{KT}\sqrt{\min\left(\sum_{i\in S_{t}}\|x_{ti}\|_{\mathbf{J}_{t}^{-1}}^{2},1\right)}.$
Hence, their regret should actually be $\tilde{\mathcal{O}}(d\sqrt{KT}+d^{2}/\kappa)$, which is worse than the result in our Theorem 2 (upper bound under uniform rewards) by a factor of $K$.